### v3

Abstract
This paper introduces deep super ReLU networks (DSRNs) as a method for approximating functions in Sobolev spaces measured by Sobolev norms $W^{m,p}$ for $m\in\mathbb{N}$ with $m\geq 2$ and $1\leq p\leq+\infty$. Standard ReLU deep neural networks (ReLU DNNs) cannot achieve this goal. DSRNs consist primarily of ReLU DNNs, and several layers of the square of ReLU added at the end to smooth the networks output. This approach retains the advantages of ReLU DNNs, leading to the straightforward training. The paper also proves the optimality of DSRNs by estimating the VC-dimension of higher-order derivatives of DNNs, and obtains the generalization error in Sobolev spaces via an estimate of the pseudo-dimension of higher-order derivatives of DNNs.
1 Introduction
Deep neural networks (DNNs) with the rectified linear unit (ReLU) activation function [13] have become increasingly popular in scientific and engineering applications, including image classification [20, 18], regularization [6], and dynamic programming [12, 39]. However, when tasked with approximating higher-order derivatives of target functions, such as in solving partial differential equations (PDEs) [21, 8, 31, 10] by DNNs, ReLU DNNs are unable to provide accurate results due to their lack of smoothness. For instance, physics-informed neural networks (PINNs) [31] have emerged as a popular approach to solving PDEs. When solving a PDE defined on a domain $\Omega=[a,b]^{d}$
$\begin{cases}\Delta u=f&\text{ in }\Omega,\\ u=g&\text{ on }\partial\Omega,\end{cases}$
the loss functions $L(\boldsymbol{\theta})$ is defined as
$L(\boldsymbol{\theta}):=\|\Delta\phi(\boldsymbol{x};\boldsymbol{\theta})-f\|^{2}_{L^{2}(\Omega)}+\beta\|\phi(\boldsymbol{x};\boldsymbol{\theta})-g\|^{2}_{L^{2}(\partial\Omega)},$
where $\phi(\boldsymbol{x};\boldsymbol{\theta})$ denotes a deep neural network (DNN) with parameters $\boldsymbol{\theta}$, and $\beta>0$ is a constant that balances the importance of the boundary and interior. However, ReLU DNNs are unsuitable for solving such PDEs because they lack the smoothness required to accurately calculate the Laplacian of $\phi(\boldsymbol{x};\boldsymbol{\theta})$, i.e., $\Delta\phi(\boldsymbol{x};\boldsymbol{\theta})$ does not exist.
In spite of the inherent smoothness limitations of ReLU-based deep neural networks, they are still widely used in training due to their simplicity, efficient computation, scale-invariance, and the fact that they are less prone to vanishing gradients. On the other hand, some smooth activation functions, such as sigmoid functions, can suffer from vanishing gradients, making them unsuitable for specific tasks. Additionally, certain smooth activation functions, such as the square of ReLU, can have complex composition formulas making training difficult.
In this paper, we propose a method for using DNNs to approximate functions in Sobolev space measured by the norms in Sobolev spaces. Specifically, the DNNs we construct in this paper use two types of activation functions, ReLU and the square of ReLU. The square of ReLU is only present in the last several layers of the DNNs, and the number of layers with the square of ReLU is proportional to the order of $\log_{2}L$, where $L$ is the depth of the DNNs. We call this kind of DNNs as deep super ReLU networks (DSRNs).
The approximation by shallow neural networks has been explored in [28, 29, 25, 40, 9, 35] and [27]. Regarding the approximation of deep neural networks (DNNs) measured by $L^{p}$-norms where $1\leq p\leq\infty$, various studies have investigated this topic, as seen in [26, 32, 33, 22, 5] and [23].The Sobolev norm approximation of DNNs has been the subject of study in several works, including [14]; [15]; [19]; [41] and [42]. In [14], ReLU DNNs were used to approximate functions measured by $W^{s,p}$ where $s\leq 1$, but their method cannot be generalized to $s>1$ due to the lack of smoothness of ReLU. Smooth activation functions were applied in [15], which achieved a non-optimal approximation rate, and training with these functions may be difficult due to the complexity of activation functions. In [19] and [42], ReLU and its squared version were utilized as DNN activation functions, yet the following three questions remain.
Firstly, the problem of approximating Sobolev functions with DNNs exhibiting super convergence rates, measured by $W^{m,p}$ norms for $p\in[1,\infty)$, remains unresolved, although the $W^{m,\infty}$ case was addressed by [42], and the case of $m=0$ was considered by [34]. Secondly, proper distributions of the squared ReLU function remain unknown. Incorporating it at any layer of a DNN with a depth of $L$ might lead to a polynomial degree of $2^{L}$ with respect to parameters, which poses challenges in training. Thirdly, the optimality of the approximation process remains unclear due to the complexity associated with higher-order derivatives of DNNs using ReLU and squared ReLU activations. To establish optimality, estimating the optimal Vapnik-Chervonenkis dimension (VC-dimension) and pseudo-dimension [2, 37, 1, 30] of their higher order derivatives, which characterize the complexity or richness of a function set, need to be estimated. Even for DNNs without derivatives, as demonstrated in [3], obtaining optimal bounds is challenging. The upper VC-dimension and pseudo-dimension for DNNs using squared ReLU is $\mathcal{O}(N^{2}L^{3}\log_{2}N\log_{2}L)$, where $L$ signifies DNN depth and $N$ denotes DNN width. However, this result is not directly applicable to proving the optimality of squared ReLU-based DNN approximation due to a polynomial gap in depth $L$.
In this paper, the deep super ReLU networks (DSRNs) is presented to address these three questions, as illustrated in Figure 1. The DSRN architecture consists of two components for a network with depth $L$: ReLU DNNs with a depth of approximately $\mathcal{O}(L)$, and the square of ReLU DNNs with a depth of approximately $\mathcal{O}(\log_{2}L)$—significantly shallower than the first component.
Figure 1: The structure of DSRNs.
The key achievement of this paper is that we demonstrate this type of DNNs can effectively approximate functions within Sobolev spaces as measured by Sobolev norms $W^{m,p}$ for $m\geq 2$. This result offers benefits both in theory and training practice. From a theoretical perspective, the proposed approach yields a nearly optimal approximation rate concerning the width $N$ and depth $L$ of DSRNs. This is made possible by the fact that the square of ReLU only appears in a shallower layer, approximately $\mathcal{O}(\log_{2}L)$, thereby maintaining VC-dimension and pseudo-dimension within $\mathcal{O}(N^{2}L^{2}\log_{2}N\log_{2}L)$ and eliminating the polynomial gap in the squared ReLU-based DNNs. In [42], the optimality of deep neural networks was proven only for networks where all activation functions are ReLU. The inclusion of mixed activation functions increased the complexity of the neural network space, creating a gap between space complexity and the approximation rate due to the frequent occurrence of the square of ReLU.
In terms of training, our DSRNs predominantly employ ReLU activations, with the square of ReLU appearing only at the end of the DNNs. This enhances the smoothness of DNNs while preserving a straightforward network structure without intricate composition formulas. This contrasts with [42], which did not specify where to apply the square of ReLU, leading to training difficulties. Our experiments in Sec. 7 demonstrate that our proposed DNN (DSRN) trains effectively, unlike those using the square of ReLU throughout, which are difficult to train due to gradient vanishing.
Finally, we estimate the generalization error of DSRNs using two methods. One method (Theorem 6.1) provides a better rate with respect to the number of parameters, while the other method (Theorem 6.2) achieves a better rate concerning the number of sample points but has a relative worse rate regarding the number of parameters. Both methods apply estimates of the VC-dimension and pseudo-dimension of higher-order derivatives of deep neural networks. To our knowledge, this is the first work to consider the VC-dimension and pseudo-dimension of higher-order derivatives in neural networks, which can be directly applied in the error estimation of Sobolev training.
Our main contributions are:
$\bullet$ We introduce deep super ReLU networks for approximating functions in Sobolev spaces, measured by norms in $W^{m,p}$ for $m\geq 2$. In our DNNs, ReLU is the predominant activation function, with the square of ReLU appearing only in the final layers. This approach enhances the smoothness of the DNNs while maintaining a simple network structure without intricate composition formulas.
$\bullet$ We estimate the upper bound of the VC-dimension of higher-order derivatives of DNNs. By utilizing our estimation of the VC-dimension, we demonstrate the optimality of our DNN approximation, as measured by Sobolev norms.
$\bullet$ We provide an upper bound estimation for the pseudo-dimension of high-order derivatives of DNNs. By utilizing this estimation, we can obtain the generalization error of DNNs in Sobolev spaces that contain high-order derivatives of DNNs.
2 Preliminaries
2.1 Neural networks
Let us summarize all basic notations used in the DNNs as follows:
1. Matrices are denoted by bold uppercase letters. For example, $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ is a real matrix of size $m\times n$ and $\boldsymbol{A}^{\mathrm{T}}$ denotes the transpose of $\boldsymbol{A}$.
2. Vectors are denoted by bold lowercase letters. For example, $\boldsymbol{v}\in\mathbb{R}^{n}$ is a column vector of size $n$. Furthermore, denote $\boldsymbol{v}(i)$ as the $i$-th elements of $\boldsymbol{v}$.
3. For a $d$-dimensional multi-index $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\cdots\alpha_{d}]\in\mathbb{N}^{d}$, we denote several related notations as follows:
$(a)~{}|\boldsymbol{\alpha}|=\left|\alpha_{1}\right|+\left|\alpha_{2}\right|+\cdots+\left|\alpha_{d}\right|;$
$(b)~{}\boldsymbol{x}^{\boldsymbol{\alpha}}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{d}^{\alpha_{d}},~{}\boldsymbol{x}=\left[x_{1},x_{2},\cdots,x_{d}\right]^{\mathrm{T}};$
$(c)~{}\boldsymbol{\alpha}!=\alpha_{1}!\alpha_{2}!\cdots\alpha_{d}!.$
4. Assume $\boldsymbol{n}\in\mathbb{N}_{+}^{m}$, and $f$ and $g$ are functions defined on $\mathbb{N}_{+}^{m}$, then $f(\boldsymbol{n})=\mathcal{O}(g(\boldsymbol{n}))$ means that there exists positive $C$ independent of $\boldsymbol{n},f,g$ such that $f(\boldsymbol{n})\leq Cg(\boldsymbol{n})$ when all entries of $\boldsymbol{n}$ go to +$\infty$+ ∞.
5. Define $\sigma_{1}(x):=\sigma(x)=\max\{0,x\}$ and $\sigma_{2}:=\sigma^{2}(x)$. We call the neural networks with activation function $\sigma_{t}$ with $t=i$ as $\sigma_{i}$ neural networks ($\sigma_{i}$-NNs), $i=1,2$. With the abuse of notations, we define $\sigma_{i}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ as $\sigma_{i}(\boldsymbol{x})=\left[\begin{array}[]{c}\sigma_{i}(x_{1})\\ \vdots\\ \sigma_{i}(x_{d})\end{array}\right]$ for any $\boldsymbol{x}=\left[x_{1},\cdots,x_{d}\right]^{T}\in\mathbb{R}^{d}$.
6. Define $L,N\in\mathbb{N}_{+}$, $N_{0}=d$ and $N_{L+1}=1$, $N_{i}\in\mathbb{N}_{+}$ for $i=1,2,\ldots,L$, then a $\sigma_{i}$-NN $\phi$ with the width $N$ and depth $L$ can be described as follows:
$\boldsymbol{x}=\tilde{\boldsymbol{h}}_{0}\stackrel{{\scriptstyle W_{1},b_{1}}}{{\longrightarrow}}\boldsymbol{h}_{1}\stackrel{{\scriptstyle\sigma_{i}}}{{\longrightarrow}}\tilde{\boldsymbol{h}}_{1}\ldots\stackrel{{\scriptstyle W_{L},b_{L}}}{{\longrightarrow}}\boldsymbol{h}_{L}\stackrel{{\scriptstyle\sigma_{i}}}{{\longrightarrow}}\tilde{\boldsymbol{h}}_{L}\stackrel{{\scriptstyle W_{L+1},b_{L+1}}}{{\longrightarrow}}\phi(\boldsymbol{x})=\boldsymbol{h}_{L+1},$
where $\boldsymbol{W}_{i}\in\mathbb{R}^{N_{i}\times N_{i-1}}$ and $\boldsymbol{b}_{i}\in\mathbb{R}^{N_{i}}$ are the weight matrix and the bias vector in the $i$-th linear transform in $\phi$, respectively, i.e., $\boldsymbol{h}_{i}:=\boldsymbol{W}_{i}\tilde{\boldsymbol{h}}_{i-1}+\boldsymbol{b}_{i},~{}\text{ for }i=1,\ldots,L+1$ and $\tilde{\boldsymbol{h}}_{i}=\sigma_{i}\left(\boldsymbol{h}_{i}\right),\text{ for }i=1,\ldots,L.$ In this paper, an DNN with the width $N$ and depth $L$, means (a) The maximum width of this DNN for all hidden layers is less than or equal to $N$. (b) The number of hidden layers of this DNN is less than or equal to $L$.
2.2 Sobolev spaces
Denote $\Omega$ as $[a,b]^{d}$, $D$ as the weak derivative of a single variable function, $D^{\boldsymbol{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$ as the partial derivative of a multivariable function, where $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the derivative in the $i$-th variable.
Definition 2.1 (Sobolev Spaces [11]).
Let $n\in\mathbb{N}$ and $1\leq p\leq\infty$. Then we define Sobolev spaces
$W^{n,p}(\Omega):=\left\{f\in L^{p}(\Omega):D^{\boldsymbol{\alpha}}f\in L^{p}(\Omega)\text{ for all }\boldsymbol{\alpha}\in\mathbb{N}^{d}\text{ with }|\boldsymbol{\alpha}|\leq n\right\}$
with a norm $\|f\|_{W^{n,p}(\Omega)}:=\left(\sum_{0\leq|\alpha|\leq n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{p}(\Omega)}^{p}\right)^{1/p}$, if $p<\infty$, and
$\|f\|_{W^{n,\infty}(\Omega)}:=\max_{0\leq|\alpha|\leq n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{\infty}(\Omega)}.$
Furthermore, for $\boldsymbol{f}=(f_{1},f_{2},\ldots,f_{d})$, $\boldsymbol{f}\in W^{1,\infty}(\Omega,\mathbb{R}^{d})$ if and only if $f_{i}\in W^{1,\infty}(\Omega)$ for each $i=1,2,\ldots,d$ and
$\|\boldsymbol{f}\|_{W^{1,\infty}(\Omega,\mathbb{R}^{d})}:=\max_{i=1,\ldots,d}\{\|f_{i}\|_{W^{1,\infty}(\Omega)}\}.$
Definition 2.2 (Sobolev semi-norm [11]).
Let $n\in\mathbb{N}_{+}$ and $1\leq p\leq\infty$. Then we define Sobolev semi-norm $|f|_{W^{n,p}(\Omega)}:=\left(\sum_{|\alpha|=n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{p}(\Omega)}^{p}\right)^{1/p}$, if $p<\infty$, and $|f|_{W^{n,\infty}(\Omega)}:=\max_{|\alpha|=n}\left\|D^{\boldsymbol{\alpha}}f\right\|_{L^{\infty}(\Omega)}$. Furthermore, for $\boldsymbol{f}\in W^{1,\infty}(\Omega,\mathbb{R}^{d})$, we define
$|\boldsymbol{f}|_{W^{1,\infty}(\Omega,\mathbb{R}^{d})}:=\max_{i=1,\ldots,d}\{|f_{i}|_{W^{1,\infty}(\Omega)}\}.$
3 Deep Super ReLU Networks
This paper focuses on deep neural networks, which are characterized by a large value of $L$. It is well-known that ReLU neural networks are unable to approximate functions in the $W^{m,\infty}$ norm for $m\geq 2$. An innovative neural network architecture is introduced in [42], which cleverly incorporates both ReLU and squared ReLU activation functions. This architecture proves to be highly effective in the approximation of functions within Sobolev spaces, as evidenced by their alignment with Sobolev norms $W^{m,\infty}((0,1)^{d})$ for $m\geq 2$:
Corollary 3.1 ([42]).
For any $f\in W^{n,\infty}((0,1)^{d})$ with $\|f\|_{W^{n,\infty}((0,1)^{d})}\leq 1$, any $N,L,m\in\mathbb{N}_{+}$ with $NL+2^{\left\lfloor\log_{2}N\right\rfloor}\geq\max\{d,n\}$ and $L\geq\left\lceil\log_{2}N\right\rceil$, there is a $\sigma_{2}$-NN $\varphi(\boldsymbol{x})$ with the width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
$\|f(\boldsymbol{x})-\varphi(\boldsymbol{x})\|_{W^{m,\infty}((0,1)^{d})}\leq C_{11}(n,d,m)N^{-2(n-m)/d}L^{-2(n-m)/d},$
where $C_{11}$ is the constant independent of $N,L$.
Nevertheless, a limitation of this corollary that uses ReLU and its square as activation functions in DNNs is that the distribution of the square of ReLU is yet to be determined. If the distribution is entirely unknown, training complexity may significantly increase. Hence, this paper introduces a novel DNN architecture that predominantly uses ReLU activation functions, allowing the square of ReLU to appear only in the last several layers. This design allows us to achieve a nearly optimal rate of approximation, reduce the complexity of DNNs as stated in Corollary 3.1, and obtain a better generalization error, making the training process easier. We will rigorously discuss this structure later and show by the numerical experiment that this network performs better in training compared to networks where the activation functions are either all ReLU or all square of ReLU in Sec. 7.
We define a subset of $\sigma_{2}$-NNs with $L\gg 1$ and $C=\mathcal{O}(1)$ with respect to $L$ as follows:
$\mathcal{N}_{C,L}:=$$\{\psi(\boldsymbol{x})=\psi_{2}\circ\boldsymbol{\psi}_{1}(\boldsymbol{x}):\psi_{2}\text{ is a $\sigma_{2}$-NN with depth $L_{2}$,}$
[…]
We refer to elements in $\mathcal{N}_{C,L}$ as deep super ReLU networks (DSRNs).
4 Approximation by DSRNs for Functions in Sobolev Spaces
In this section, we investigate the ability of deep super ReLU networks (DSRNs) to approximate functions in Sobolev spaces. Specifically, we consider the Sobolev space $W^{m,p}([a,b]^{d})$ for $m\geq 2$. We establish that DSRNs can effectively approximate functions in $W^{m,p}([a,b]^{d})$ with a nearly optimal rate of convergence. Without loss of generality, we consider the case of the $m=2$.
Theorem 4.1.
For any $f\in W^{n,p}([a,b]^{d})$ for $p\in[1,\infty]$, any $d,n\geq 2$ and $N,L\in\mathbb{N}_{+}$ with $\log N\geq d(\log n+\log d)$, there is a DSRN $\gamma(\boldsymbol{x})$ in $\mathcal{N}_{1,9(L+1)\log_{2}4L}$ with the width $2^{d+6}n^{d+1}(N+d)\log_{2}(8N)$ such that
$\|f(\boldsymbol{x})-\gamma(\boldsymbol{x})\|_{W^{2,p}([a,b]^{d})}\leq 2^{d+7}C_{10}(n,d)\|f(\boldsymbol{x})\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
where $C_{10}$ is the constant independent of $N,L$.
Remark 4.2.
As we can see, the approximation rate of the DSRN in Theorem 4.1 is
$\mathcal{O}(N^{-2(n-2)/d}L^{-2(n-2)/d})$
with $\mathcal{O}(N^{2}L)$ parameters, which is better than traditional methods and achieves super-convergence. This rate is achieved based on the bit extraction method, where the super-convergence approximation error decreases faster than the metric entropy of the space of Sobolev functions. Therefore, observing this error order in numerical experiments can be challenging, as discuss the details in [34].
4.1 Proof Sketch for Theorem 4.1
The proof of Theorem 4.1 can be outlined in five parts, and the complete proof is provided in next subsections:
(i): First, define a sequence of subsets of $\Omega$ in Definition 4.5.
Then we define a partition of unity $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ on $[a,b]^{d}$ with supp $\text{supp }g_{\boldsymbol{m}}\cap[a,b]^{d}\subset\Omega_{\boldsymbol{m}}$ for each $\boldsymbol{m}\in\{1,2\}^{d}$ in Definition 4.7.
(ii): Next we use the Proposition 4.8 to present $\{s_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ by $\sigma_{2}$-NNs.
(iii): Third, for each $\Omega_{\boldsymbol{m}}\subset[a,b]^{d}$, where $\boldsymbol{m}\in\{1,2\}^{d}$, we find a function $f_{K,\boldsymbol{m}}$ satisfying
$\|f-f_{K,\boldsymbol{m}}\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-2)},$
$\|f-f_{K,\boldsymbol{m}}\|_{W^{1,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-1)},$
(2) $\|f-f_{K,\boldsymbol{m}}\|_{L^{p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-n},$
where $C_{1}$ is a constant independent of $K$. Moreover, each $f_{K,\boldsymbol{m}}$ can be expressed as $f_{K,\boldsymbol{m}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}$, where $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})$ is a piecewise constant function on $\Omega_{\boldsymbol{m}}$. The proof of this result is based on the Bramble-Hilbert Lemma [4, Lemma 4.3.8].
(iv): The fourth step involves approximating $f_{K,\boldsymbol{m}}$ using DSRN $\gamma_{\boldsymbol{m}}$, following the approach outlined in [26]. This method is suitable for our work because $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})$ is a piecewise constant function on $\Omega_{\boldsymbol{m}}$, and the weak derivative of $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})$ on $\Omega_{\boldsymbol{m}}$ is zero. This property allows for the use of the $L^{\infty}$ norm approximation method presented in [26]. Thus, we obtain a neural network $\psi_{\boldsymbol{m}}$ with width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
(3) $\|f_{K,\boldsymbol{m}}-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
where $C$ is a constant independent of $N$ and $L$.
By combining (iii) and (iv) and setting $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, we obtain that for each $\boldsymbol{m}\in\{1,2\}^{d}$, there exists a DSRN $\gamma_{\boldsymbol{m}}$ with width $\mathcal{O}(N\log N)$ and depth $\mathcal{O}(L\log L)$ such that
(4) $\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{10}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d},$
where $C_{10}$ is a constant independent of $N$ and $L$.
(v): The final step is to combine the sequences $\{s_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ and $\{\gamma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ to construct a network that can approximate $f$ over the entire space $[a,b]^{d}$. We define the sequence $\{s_{\boldsymbol{m}}\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ because $\gamma_{\boldsymbol{m}}$ may not accurately approximate $f$ on $[a,b]^{d}\backslash\Omega_{\boldsymbol{m}}$. The purpose of $s_{\boldsymbol{m}}$ is to remove this portion of the domain and to allow other networks to approximate $f$ on $[a,b]^{d}\backslash\Omega_{\boldsymbol{m}}$.
4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
First, we list a few basic lemmas of $\sigma_{i}$ neural networks repeatedly applied in our main analysis.
Lemma 4.3.
The following basic lemmas of $\sigma_{2}$ neural networks should hold:
(i) $f(x)=x^{2}$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and two neurons, and $f(x,y)=xy=\frac{(x+y)^{2}-(x-y)^{2}}{4}$ can be realized exactly by a $\sigma_{2}$ neural network with one hidden layer and four neurons.
(ii) Let $P(\boldsymbol{x})$ be a polynomial of total degree $m$ in $\mathbb{R}^{d}$. There is a $\sigma_{2}$-NN of depth $\lceil\log_{2}m\rceil$ and width $2\,(m+1)^{d}$ that exactly represents the polynomial $P(\boldsymbol{x})$.
(iii) For any $s>0$, there is a $\sigma_{2}$-NN with width $2(s+1)$ and depth $\log_{2}m$ that exactly represents $\text{ReLU}^{s+2}$.
Proof 4.4.
(i) is easy to check by $x^{2}=\sigma_{2}(x)+\sigma(-x)$. (ii) is the results in [17]. For (iii), we know that
(5) $\text{ReLU}^{s+2}(x)=x^{s}\cdot\text{ReLU}^{2}(x).$
Then combining (i) and (ii), we have that $\text{ReLU}^{s+2}$ is a $\sigma_{2}$ neural network with width $2(s+1)$ and depth $\log_{2}m$.
Define a sequence of subsets of $\Omega$.
Definition 4.5.
Given $K,d\in\mathbb{N}^{+}$, and for any $\boldsymbol{m}=(m_{1},m_{2},\ldots,m_{d})\in\{1,2\}^{d}$, we define $\Omega_{\boldsymbol{m}}:=\prod_{j=1}^{d}\Omega_{m_{j}},$ where
$\Omega_{1}$ $:=\bigcup_{i=0}^{K-1}\left[a+\frac{i\cdot(b-a)}{K},a+\frac{i\cdot(b-a)}{K}+\frac{3\cdot(b-a)}{4K}\right],$
(6) $\Omega_{2}$ $:=\bigcup_{i=0}^{K}\left[\frac{i\cdot(b-a)}{K}-\frac{b-a}{2K},a+\frac{i\cdot(b-a)}{K}+\frac{b-a}{4K}\right]\cap[a,b].$
Next, we define a function which will be repeatly used in the proof in this section.
Definition 4.6.
Define $s(x)$ from $\mathbb{R}\to[0,1]$ as
(7) $s(x):=\begin{cases}2x^{2},~{}&x\in\left[0,\frac{1}{2}\right]\\ -2(x-1)^{2}+1,~{}&x\in\left[\frac{1}{2},1\right]\\ 1,~{}&x\in\left[1,2\right]\\ -2(x-2)^{2}+1,~{}&x\in\left[2,\frac{5}{2}\right]\\ 2(x-3)^{2},~{}&x\in\left[\frac{5}{2},3\right]\\ 0,~{}&\text{otherwise}.\end{cases}$
Figure 2: $s(x)$ in $\mathbb{R}$.
Definition 4.7.
Given $K\in\mathbb{N}_{+}$, we define two functions in $\mathbb{R}$:
(8) $s_{1}(x)=\sum_{i=0}^{K}s\left(4Kx+1-4i\right),~{}s_{2}(x)=s_{1}\left(x+\frac{1}{2K}\right).$
Then for any $\boldsymbol{m}=(m_{1},m_{2},\ldots,m_{d})\in\{1,2\}^{d}$, we define
(9) $s_{\boldsymbol{m}}(\boldsymbol{x}):=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)$
for any $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{d})\in\mathbb{R}^{d}$.
Proposition 4.8.
Given $N,L,d\in\mathbb{N}_{+}$ with $\log_{2}N>d\log d$, and the setting $K=\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor$, $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ defined in Definition 4.7 satisfies:
(i): $\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{L^{\infty}([a,b]^{d})}\leq 1$$\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{1,\infty}([a,b]^{d})}\leq\frac{8K}{b-a}$ and $\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,\infty}([a,b]^{d})}\leq\frac{64K^{2}}{(b-a)^{2}}$ for any $\boldsymbol{m}\in\{1,2\}^{d}$.
(ii): $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ is a partition of the unity $[a,b]^{d}$ with ${\rm supp}~{}s_{\boldsymbol{m}}(\boldsymbol{x})\cap[a,b]^{d}=\Omega_{\boldsymbol{m}}$ defined in Definition 4.5.
(iii): For any $\boldsymbol{m}\in\{1,2\}^{d}$, we have that there is a $\sigma_{1}$ neural network $\lambda_{\boldsymbol{m},1}(\boldsymbol{x})$ with the width $4N$ and depth $\lceil 2\log_{2}L\rceil$ and a $\sigma_{2}$ neural network $\lambda_{\boldsymbol{m},2}(\boldsymbol{x})$ with the width $6N$ and depth $\lceil 1+\log_{2}d\rceil$, such as
$\lambda_{\boldsymbol{m},2}\circ\lambda_{\boldsymbol{m},1}(\boldsymbol{x})=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)=s_{\boldsymbol{m}}(\boldsymbol{x}),\boldsymbol{x}\in[a,b]^{d}.$
Proof 4.9.
(i) and (ii) are proved by direct calculation. The proof of (iii) follows:
First, we architect $s(x)$ by a $\sigma_{2}$ neural network. There is a $\sigma_{1}$ neural network $g(x)$ with width 3 and one hidden layer that:
(10) $g(x):=\begin{cases}x,~{}&x\in\left[0,\frac{1}{2}\right]\\ \frac{1}{2},~{}&x\in\Big{[}\frac{1}{2},+\infty\Big{)}\\ 0,~{}&\text{otherwise}.\end{cases}$
Based on (iii) in Lemma 4.3, $g^{2}(x)$ is a $\sigma_{2}$ neural network with width 3 and two hidden layers. Then by direct calculation, we notice that
(11) $s(x)=2[x]_{+}^{2}-4[x-0.5]_{+}^{2}+2[x-1]_{+}^{2}-2[x-2]_{+}^{2}+4[x-2.5]_{+}^{2}-2[x-3]_{+}^{2}.$
which is a $\sigma_{2}$ neural network with with width 6 and one hidden layer. The $\widetilde{g}(x)$ defined as
(12) $\widetilde{g}(x)=\sum_{i=0}^{\lfloor N^{1/d}\rfloor-1}s\left(4Kx-4i-\frac{1}{2}\right)$
is a $\sigma_{2}$ neural network with the width $12(\lfloor N^{1/d}\rfloor)$ and two hidden layers.
Next, we construct $\psi_{i}$ for $i=2,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ based on the symmetry and periodicity of $\widetilde{g}(x)$ due to the symmetry of $s(x)$ (see Fig. 2). $\psi_{2}$ is the function with period $\frac{2}{\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor}$ in $\left[0,\frac{1}{\lfloor L^{2/d}\rfloor}\right]$, and each period is a hat function with gradient 1. $\psi_{i}$ for $i=3,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ is the hat function in $\left[0,\frac{2^{i-2}}{\lfloor L^{2/d}\rfloor}\right]$. The schematic diagram is in Figure 3 (The diagram is shown the case for $\lfloor N^{1/d}\rfloor$ is a even integer).
Figure 3: $\psi_{i}$ for $i=2,\ldots,\lceil\frac{1}{d}\log_{2}L\rceil$.
Note that $\psi_{2}\circ\psi_{3}\ldots\circ\psi_{\lceil\frac{2}{d}\log_{2}L\rceil+2}(x)$ is the function with period $\frac{2}{\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor}$ in $[0,1]\subset\left[0,\frac{2^{\lceil\frac{2}{d}\log_{2}L\rceil}}{\lfloor L^{2/d}\rfloor}\right]$, and each period is a hat function with gradient 1. Then function
(13) $\psi(x):=\widetilde{g}(x)\circ\psi_{2}\circ\psi_{3}\ldots\circ\psi_{\lceil\frac{2}{d}\log_{2}L\rceil+2}(x):=\widetilde{g}\circ\psi_{1}$
is obtained by repeating reflection $\widetilde{g}(x)$ in $[0,1]$.
$\psi_{2}$ is a network with width $4\lfloor N^{1/d}\rfloor$ and one hidden layer. $\psi_{i}$ for $i=3,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ is a $\sigma_{1}$-NN with width 3 and one hidden layer. Hence $\psi_{1}(x)$ is a $\sigma_{1}$-NN with width $4\lfloor N^{1/d}\rfloor$ and depth $\lceil\frac{2}{d}\log_{2}L\rceil$ and $s_{1}(x)=\psi\left(x+\frac{1}{8K}\right)$ and $s_{2}(x)=\psi\left(x+\frac{5}{8K}\right)$.
Based on (ii) in Lemma 4.3, we have that there is a $\sigma_{1}$ neural network $\lambda_{\boldsymbol{m},1}(\boldsymbol{x})$ with the width $4N$ and depth $\lceil 2\log_{2}L\rceil$ and a $\sigma_{2}$ neural network $\lambda_{\boldsymbol{m},2}(\boldsymbol{x})$ with the width $6N$ and depth $\lceil 1+\log_{2}d\rceil$, such as
$\lambda_{\boldsymbol{m},2}\circ\lambda_{\boldsymbol{m},1}(\boldsymbol{x})=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)=s_{\boldsymbol{m}}(\boldsymbol{x}),\boldsymbol{x}\in[a,b]^{d},$
where the order of width is based on $dN^{1/d}\leq N$ for $N>d$.
Next, we present some propositions for ReLU DNNs.
Proposition 4.10 ([26]).
Given any $N,L\in\mathbb{N}_{+}$ and $\delta\in\Big{(}0,\frac{1}{3K}\Big{]}$ for $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, there exists a $\sigma_{1}$-NN $\phi$ with the width $54N+$ 5 and depth $44L+$ 4 such that
$\phi(x)=k,x\in\left[a+\frac{k\cdot(b-a)}{K},a+\frac{(k+1)\cdot(b-a)}{K}-\delta\cdot 1_{k<a+(K-1)\cdot(b-a)}\right],~{}k=0,\ldots,K-1.$
Remark 4.11.
In [26], this proposition is proved on the $[0,1]$ domain. Here, we extend it to the $[a,b]$ domain, which is straightforward by applying a rescaling and transformation.
Proposition 4.12 ([26]).
Given any $N,L,s\in\mathbb{N}_{+}$ and $\xi_{i}\in[0,1]$ for $i=0,1,\ldots N^{2}L^{2}-1$, there exists a $\sigma_{1}$-NN $\phi$ with the width $16s(N+1)\log_{2}(8N)$ and depth $(5L+2)\log_{2}(4L)$ such that
1. $|\phi(i)-\xi_{i}|\leq N^{-2s}L^{-2s}$ for $i=0,1,\ldots N^{2}L^{2}-1$.
2. $0\leq\phi(x)\leq 1$, $x\in\mathbb{R}$.
4.3 Proof of Theorem 4.1
To derive the approximation result, it can be broken down into three sequential steps. Without sacrificing generality, we focus on approximating functions in $W^{n,p}$ using the norm measured in $W^{m,p}$.
4.3.1 Function approximation in Sobolev spaces through the Bramble–Hilbert Lemma [4, Lemma 4.3.8]
Before we show Proposition 4.13, we define subsets of $\Omega_{\boldsymbol{m}}$ for simplicity notations.
For any $\boldsymbol{m}\in\{1,2\}^{d}$, we define
(14) $\Omega_{\boldsymbol{m},\boldsymbol{i}}:=[a,b]^{d}\cap\prod_{j=1}^{d}\left[a+\frac{2i_{j}-1_{m_{j}<2}}{2K}\cdot(b-a),a+\frac{3+4i_{j}-2\cdot 1_{m_{j}<2}}{4K}\cdot(b-a)\right]$
$\boldsymbol{i}=(i_{1},i_{2},\ldots,i_{d})\in\{0,1\ldots,K\}^{d}$, and it is easy to check $\bigcup_{\boldsymbol{i}\in\{0,1\ldots,K\}^{d}}\Omega_{\boldsymbol{m},\boldsymbol{i}}=\Omega_{\boldsymbol{m}}$.
Proposition 4.13.
Let $K\in\mathbb{N}_{+}$, $p\in[1,\infty]$ and $n\geq 2$. Then for any $f\in W^{n,p}([a,b]^{d})$ and $\boldsymbol{m}\in\{1,2\}^{d}$, there exist piece-wise polynomials function $f_{K,\boldsymbol{m}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}$ on $\Omega_{\boldsymbol{m}}$ (Definition 4.5) with the following properties:
$\|f-f_{K,\boldsymbol{m}}\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-2)},$
$\|f-f_{K,\boldsymbol{m}}\|_{W^{,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-1)},$
(15) $\|f-f_{K,\boldsymbol{m}}\|_{L^{p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-n}.$
Furthermore, $g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x}):\Omega_{\boldsymbol{m}}\to\mathbb{R}$ is a constant function with on each $\Omega_{\boldsymbol{m},\boldsymbol{i}}$ for $\boldsymbol{i}\in\{0\ldots,K\}^{d}.$ And
(16) $|g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})|\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}})}(4K)^{\frac{d}{p}}$
for all $\boldsymbol{x}\in\Omega_{\boldsymbol{m}}$, where $C_{1}$ and $C_{2}$ are constants independent of $K$.
This proof is similar to that of [14, Lemma C.4.], but we provide detailed proof as follows for readability. Before the proof, we must introduce the partition of unity, average Taylor polynomials, and a lemma.
Definition 4.14 (The partition of unity).
Let $d,K\in\mathbb{N}_{+}$, then
$\Psi=\left\{h_{\boldsymbol{i}}:{\boldsymbol{i}}\in\{0,1,\ldots,K\}^{d}\right\}$
with $h_{\boldsymbol{i}}:\mathbb{R}^{d}\to\mathbb{R}$ for all ${\boldsymbol{i}}\in\{0,1,\ldots,K\}^{d}$ is called the partition of unity $[a,b]^{d}$ if it satisfies
(i): $0\leq h_{\boldsymbol{i}}(\boldsymbol{x})\leq 1$ for every $h_{\boldsymbol{i}}\in\Psi$.
(ii): $\sum_{h_{\boldsymbol{i}}\in\Psi}h_{\boldsymbol{i}}=1$ for every $x\in[a,b]^{d}$.
Definition 4.15.
Let $n\geq 1$ and $f\in W^{n,p}([a,b]^{d})$, $\boldsymbol{x}_{0}\in[a,b]^{d}$ and $r>0$ such that for the ball $B(\boldsymbol{x}_{0}):=B(\boldsymbol{x}_{0})_{r,|\cdot|}$ which is a compact subset of $[a,b]^{d}$. The corresponding Taylor polynomial of order $n$ of $f$ averaged over $B$ is defined for
(17) $Q^{n}f(x):=\int_{B}T^{n}_{\boldsymbol{y}}f(\boldsymbol{x})b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}$
where
$T^{n}_{\boldsymbol{y}}{f}(\boldsymbol{x})$ $:=\sum_{|\boldsymbol{\alpha}|\leq n-1}\frac{1}{\boldsymbol{\alpha}!}D^{\boldsymbol{\alpha}}{f}(\boldsymbol{y})(\boldsymbol{x}-\boldsymbol{y})^{\boldsymbol{\alpha}},$
$b_{r}(\boldsymbol{x})$ :=$:=\begin{cases}\frac{1}{c_{r}}e^{-\left(1-\left(|\boldsymbol{x}-\boldsymbol{x}_{0}|/r\right)^{2}\right)^{-1}},~{}&|\boldsymbol{x}-\boldsymbol{x}_{0}|<r,\\ 0,~{}&|\boldsymbol{x}-\boldsymbol{x}_{0}|\leq r,\end{cases}$
(18) $c_{r}$ $=\int_{\mathbb{R}^{d}}e^{-\left(1-\left(|\boldsymbol{x}-\boldsymbol{x}_{0}|/r\right)^{2}\right)^{-1}}\,\mathrm{d}x.$
Lemma 4.16 ([14]).
Let $n\geq 1$ and $f\in W^{n,p}([a,b]^{d})$, $\boldsymbol{x}_{0}\in\Omega$ and $r>0$ such that for the ball $B:=B_{r,|\cdot|}(\boldsymbol{x}_{0})$ which is a compact subset of $[a,b]^{d}$. The corresponding Taylor polynomial of order $n$ of $f$ averaged over $B$ can be read as
$Q^{n}f(\boldsymbol{x})=\sum_{|\boldsymbol{\alpha}|\leq n-1}c_{f,\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}}.$
Furthermore,
(19) $\left|c_{f,\boldsymbol{\alpha}}\right|\leq C_{2}(n,d)r^{-\frac{d}{p}}\|{f}\|_{W^{n-1,p}(B)}$
where $C_{2}(n,d)$ is a constant independent of $f$.
Proof 4.17.
Based on [14, Lemma B.9.], $Q^{n}f(x)$ can be read as
(20) $Q^{n}f(\boldsymbol{x})=\sum_{|\boldsymbol{\alpha}|\leq n-1}c_{f,\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}}$
where
(21) $c_{f,\boldsymbol{\alpha}}=\sum_{|\boldsymbol{\alpha}+\boldsymbol{\beta}|\leq n-1}\frac{1}{(\boldsymbol{\beta}+\boldsymbol{\alpha})!}a_{\boldsymbol{\beta}+\boldsymbol{\alpha}}\int_{B}D^{\boldsymbol{\alpha}+\boldsymbol{\beta}}{f}(\boldsymbol{x})\boldsymbol{y}^{\boldsymbol{\beta}}b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}$
for $a_{\boldsymbol{\beta}+\boldsymbol{\alpha}}\leq\frac{(\boldsymbol{\alpha}+\boldsymbol{\beta})!}{\boldsymbol{\alpha}!\boldsymbol{\beta}!}$. Note that
$\left|\int_{B}D^{\boldsymbol{\alpha}+\boldsymbol{\beta}}{f}(\boldsymbol{x})\boldsymbol{y}^{\boldsymbol{\beta}}b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}\right|\leq\left|\int_{B}D^{\boldsymbol{\alpha}+\boldsymbol{\beta}}{f}(\boldsymbol{x})b_{r}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}\right|$
(22) $\leq$ $\|{f}\|_{W^{n-1,p}(B)}\|b_{r}(x)\|_{L^{q}(B)}\leq cr^{-\frac{d}{p}}\|{f}\|_{W^{n-1,p}(B)}.$
Then
(23) $\left|c_{f,\boldsymbol{\alpha}}\right|\leq C_{2}(n,d)r^{-\frac{d}{p}}\|{f}\|_{W^{n-1,p}(B)}$
where $C_{2}(n,d)=\sum_{|\boldsymbol{\alpha}+\boldsymbol{\beta}|\leq n-1}\frac{c}{\boldsymbol{\alpha}!\boldsymbol{\beta}!}$.
The proof of Proposition 4.13 is based on average Taylor polynomials and the Bramble–Hilbert Lemma [4, Lemma 4.3.8].
Definition 4.18.
Let $\Omega,~{}B\in\mathbb{R}^{d}$. Then $\Omega$ is called stared-shaped with respect to $B$ if
$\overline{\text{conv}}\left(\{\boldsymbol{x}\}\cup B\subset\Omega\right),~{}\text{for all }\boldsymbol{x}\in\Omega.$
Definition 4.19.
Let $\Omega\in\mathbb{R}^{d}$ be bounded, and define
:=$\mathcal{R}:=\left\{r>0:\begin{array}[]{l}\text{ there exists }\boldsymbol{x}_{0}\in\Omega\text{ such that }\Omega\text{ is }\\ \text{ star-shaped with respect to }B_{r,|\cdot|}\left(\boldsymbol{x}_{0}\right)\end{array}\right\}.$
Then we define
$r_{\max}^{\star}:=\sup\mathcal{R}\quad\text{ and call }\quad\gamma:=\frac{\operatorname{diam}(\Omega)}{r_{\max}^{\star}}$
the chunkiness parameter of $\Omega$ if $\mathcal{R}\not=\emptyset$.
Lemma 4.20 (Bramble–Hilbert Lemma [4, Lemma 4.3.8]).
Let $\Omega\in\mathbb{R}^{d}$ be open and bounded, $\boldsymbol{x}_{0}\in\Omega$ and $r>0$ such that $\Omega$ is the stared-shaped with respect to $B:=B_{r,|\cdot|}\left(\boldsymbol{x}_{0}\right)$, and $r\geq\frac{1}{2}r_{\max}^{\star}$. Moreover, let $n\in\mathbb{N}_{+}$, $1\leq p\leq p$ and denote by $\gamma$ by the chunkiness parameter of $\Omega$. Then there is a constant $C(n,d,\gamma)>0$ such that for all $f\in W^{n,p}(\Omega)$
$\left|f-Q^{n}f\right|_{W^{k,p}(\Omega)}\leq C(n,d,\gamma)h^{n-k}|f|_{W^{n,p}(\Omega)}\quad\text{ for }k=0,1,\ldots,n$
where $Q^{n}f$ denotes the Taylor polynomial of order $n$ of $f$ averaged over $B$ and $h=\operatorname{diam}(\Omega)$.
Proof 4.21 (Proof of Proposition 4.13).
Without loss of generalization, we prove the case for $\boldsymbol{m}=(1,1,\ldots,1)=:\boldsymbol{m}_{*}$.
Denote $E:W^{n,p}([a,b]^{d})\to W^{n,p}(\mathbb{R}^{d})$ be an extension operator [36] and set $\tilde{f}:=Ef$ and $C_{E}$ is the norm of the extension operator.
Define $p_{f,\boldsymbol{i}}$ as the average Taylor polynomial Definition 4.15 in
$B_{\boldsymbol{i},K}:=B_{\frac{1}{4K}\cdot(b-a),|\cdot|}\left(a+\frac{8\boldsymbol{i}+3}{8K}\cdot(b-a)\right)$
i.e.
(24) $p_{f,\boldsymbol{i}}:=\int_{B_{\boldsymbol{i},K}}T^{n}_{\boldsymbol{y}}\tilde{f}(\boldsymbol{x})b_{\frac{1}{4K}}(\boldsymbol{y})\,\mathrm{d}\boldsymbol{y}.$
Based on Lemma 4.16, $p_{f,\boldsymbol{i}}$ can be read as
(25) $p_{f,\boldsymbol{i}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}c_{f,\boldsymbol{i},\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}},$
where
(26) $|c_{f,\boldsymbol{i},\boldsymbol{\alpha}}|\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(B_{\boldsymbol{i},K})}(4K)^{\frac{d}{p}}\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}.$
The reason to define average Taylor polynomial on $B_{\boldsymbol{i},K}$ is to use the Bramble–Hilbert Lemma 4.20 on
$\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}=B_{\frac{3}{8K}\cdot(b-a),\|\cdot\|_{\ell_{\infty}}}\left(a+\frac{8\boldsymbol{i}+3}{8K}\cdot(b-a)\right)=\prod_{j=1}^{d}\left[a+\frac{i_{j}}{K}\cdot(b-a),a+\frac{3+4i_{j}}{4K}\cdot(b-a)\right].$
Note that
$\frac{b-a}{4K}\geq\frac{b-a}{2}\cdot\frac{3}{8K}=\frac{1}{2}r_{\max}^{\star}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}),~{}\gamma(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})=\frac{\operatorname{diam}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}{r_{\max}^{\star}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}=2\sqrt{d}.$
Therefore we can apply the Bramble–Hilbert Lemma 4.20 and have
$\|\tilde{f}-p_{f,\boldsymbol{i}}\|_{L^{p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}$ $\leq C_{BH}(n,d)|f|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}K^{-n},$
$|\tilde{f}-p_{f,\boldsymbol{i}}|_{W^{1,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}$ $\leq C_{BH}(n,d)|f|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}K^{-(n-1)},$
(27) $|\tilde{f}-p_{f,\boldsymbol{i}}|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}$ $\leq C_{BH}(n,d)|f|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}K^{-(n-2)},$
where $C_{BH}(n,d)=|\{|\boldsymbol{\alpha}|=n\}|\frac{1}{d\int_{0}^{1}x^{d-1}e^{-\left(1-x^{2}\right)^{-1}}\,\mathrm{d}x}\left(2+4\sqrt{d}\right)^{d}C_{E}$ following the proof of Lemma [4, Lemma 4.3.8]. Therefore,
$\|\tilde{f}-p_{f,\boldsymbol{i}}\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}K^{-(n-2)},$
where $C_{1}(n,d)=3C_{BH}(n,d)$.
Now we construct a partition of unity that we use in this theorem. First of all, given any integer $K$, define $\{h_{i}\}_{i=0}^{K}$ from $\mathbb{R}\to\mathbb{R}$:
(28) $h_{i}(x):=h\left(\frac{4K}{b-a}\left(x-\frac{8i+3}{8K}-a\right)\right),\quad h(x):=\begin{cases}1,~{}&|x|<\frac{3}{2}\\ 0,~{}&|x|>2\\ 4-2|x|,~{}&\frac{3}{2}\leq|x|\leq 2.\end{cases}$
It is easy to check that $\{h_{i}\}_{i=0}^{K}$ is a partition of unity of $[a,b]$ and $h_{i}(x)=1$ for $x\in\left[a+\frac{i\cdot(b-a)}{K},a+\frac{3+4i}{4K}\cdot(b-a)\right]$. Hence we can define $h_{\boldsymbol{i}}(\boldsymbol{x})$ for $\boldsymbol{i}=(i_{1},i_{2},\ldots,i_{d})\in\{0,\ldots,K\}^{d}$ and $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{d})\in\mathbb{R}^{d}$:
(29) $h_{\boldsymbol{i}}(\boldsymbol{x})=\prod_{j=1}^{d}h_{i_{j}}(x_{j}),$
and $\left\{h_{\boldsymbol{i}}:{\boldsymbol{i}}\in\{0,1,\ldots,K\}^{d}\right\}$ is a partition of unity of $[a,b]^{d}$ and $h_{\boldsymbol{i}}(\boldsymbol{x})=1$ for
$\boldsymbol{x}\in\prod_{j=1}^{d}\left[a+\frac{i\cdot(b-a)}{K},a+\frac{3+4i}{4K}\cdot(b-a)\right]=\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}$
and $\boldsymbol{i}=(i_{1},i_{2},\ldots,i_{d})\in\{0,1,\ldots,K\}^{d}$.
Then
$\|h_{\boldsymbol{i}}(\tilde{f}-p_{f,\boldsymbol{i}})\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}K^{-(n-2)}.$
Finally, for $s=0,1,2$,
$\left\|f-\sum_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}h_{\boldsymbol{i}}p_{f,\boldsymbol{i}}\right\|^{p}_{W^{s,p}(\Omega_{m_{*}})}=$ $\sum_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}\|h_{\boldsymbol{i}}(\tilde{f}-p_{f,\boldsymbol{i}})\|^{p}_{W^{s,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}$
$\leq\left[C_{1}(n,d)K^{-(n-s)}\right]^{p}\cdot\sum_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}\|f\|^{p}_{W^{n,p}(\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}})}$
(30) $=\left[C_{1}(n,d)\|f\|^{p}_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}K^{-(n-s)}\right]^{p},$
which is due to $\cup_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}=\Omega_{m_{*}}$, $\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}\cap\Omega_{\boldsymbol{m}_{*},\boldsymbol{j}}=\emptyset$ for $\boldsymbol{i}\not=\boldsymbol{j}$ and supp $h_{\boldsymbol{i}}\cap\Omega_{m_{*}}=\Omega_{\boldsymbol{m}_{*},\boldsymbol{i}}$.
Last of all,
$f_{k,\boldsymbol{m}_{*}}(\boldsymbol{x}):=$ $\sum_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}h_{\boldsymbol{i}}p_{f,\boldsymbol{i}}=\sum_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}\sum_{|\boldsymbol{\alpha}|\leq n-1}h_{\boldsymbol{i}}c_{f,\boldsymbol{i},\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}}$
$=$ $\sum_{|\boldsymbol{\alpha}|\leq n-1}\sum_{\boldsymbol{i}\in\{0,1,\ldots,K\}^{d}}h_{\boldsymbol{i}}c_{f,\boldsymbol{i},\boldsymbol{\alpha}}\boldsymbol{x}^{\boldsymbol{\alpha}}$
(31) =:absent:=:= : $\sum_{|\boldsymbol{\alpha}|\leq n-1}g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}},$
with $|g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})|\leq C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}$ for $x\in\Omega_{\boldsymbol{m}_{*}}$. Note that $g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})$ is a step function from $\Omega_{\boldsymbol{m}_{*}}\to\mathbb{R}$:
(32) $g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})=c_{f,\boldsymbol{i},\boldsymbol{\alpha}}$
for $\boldsymbol{x}\in\prod_{j=1}^{d}\left[a+\frac{i_{j}}{K}\cdot(b-a),a+\frac{3+4i_{j}}{4K}\cdot(b-a)\right]$ and $\boldsymbol{i}=(i_{1},i_{2},\ldots,i_{d})$
4.3.2 Proof of Theorem 4.1
Proof 4.22 (Proof of Theorem 4.1).
Without loss of the generalization, we consider the case for $\boldsymbol{m}_{*}=(1,1,\ldots,1)$. Due to Proposition 4.13 and setting $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, we have
$\|f-f_{K,\boldsymbol{m}_{*}}\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}\leq C_{1}(n,d)\|f(\boldsymbol{x})\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-2)/d}L^{-2(n-2)/d}$
$\|f-f_{K,\boldsymbol{m}_{*}}\|_{W^{1,p}(\Omega_{\boldsymbol{m}_{*}})}\leq C_{1}(n,d)\|f(\boldsymbol{x})\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-1)/d}L^{-2(n-1)/d}$
(33) $\|f-f_{K,\boldsymbol{m}_{*}}\|_{L^{p}(\Omega_{\boldsymbol{m}_{*}})}\leq C_{1}(n,d)\|f(\boldsymbol{x})\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2n/d}L^{-2n/d},$
where $f_{K,\boldsymbol{m}_{*}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}$ for $x\in\Omega_{\boldsymbol{m}_{*}}$. Note that $g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})$ is a constant function for $\boldsymbol{x}\in\prod_{j=1}^{d}\left[a+\frac{i_{j}}{K}\cdot(b-a),a+\frac{3+4i_{j}}{4K}\cdot(b-a)\right]$ and $\boldsymbol{i}=(i_{1},i_{2},\ldots,i_{d})\in\{0,1,\ldots,K-1\}^{d}$. The remaining part is to approximate $f_{K,\boldsymbol{m}_{*}}$ by neural networks.
The way to approximate $g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})$ is similar with [19, Theorem 3.1]. First of all, due to Proposition 4.10, there is a neural network $\phi_{1}(x)$ with the width $54N+$ 5 and depth $44L+$ 4 such that
(34) $\phi(x)=k,x\in\left[a+\frac{k\cdot(b-a)}{K},a+\frac{(k+1)\cdot(b-a)}{K}-\frac{b-a}{4K}\right],~{}k=0,1,\ldots,K-1.$
Note that we choose $\delta=\frac{1}{4K}\leq\frac{1}{3K}$ in Proposition 4.10. Then define
$\boldsymbol{\phi}_{2}(\boldsymbol{x})=\left[\frac{\phi_{1}(x_{1})}{K},\frac{\phi_{1}(x_{2})}{K},\ldots,\frac{\phi_{1}(x_{d})}{K}\right]^{\mathrm{T}}.$
For each $p=0,1,\ldots,K^{d}-1$, there is a bijection
$\boldsymbol{\eta}(p)=[\eta_{1},\eta_{2},\ldots,\eta_{d}]\in\{0,1,\ldots,K-1\}^{d}$
such that $\sum_{j=1}^{d}\eta_{j}K^{j-1}=p$. Then define
$\xi_{\boldsymbol{\alpha},p}=\frac{g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\frac{\boldsymbol{\eta}(p)}{K}\right)+C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}}{2C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}}\in[0,1],$
where $C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}$ is the bounded of $g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}$ defined in Proposition 4.13. Therefore, based on Proposition 4.12, there is a neural network $\tilde{\phi}_{\boldsymbol{\alpha}}(x)$ with the width
$16\left(n-1+\frac{d}{p}\right)(N+1)\log_{2}(8N)$
and depth $(5L+2)\log_{2}(4L)$ such that $|\tilde{\phi}_{\boldsymbol{\alpha}}(p)-\xi_{\boldsymbol{\alpha},p}|\leq N^{-2\left(n-1+\frac{d}{p}\right)}L^{-2\left(n-1+\frac{d}{p}\right)}$ for $p=0,1,\ldots K^{d}-1$. Denote
$\phi_{\boldsymbol{\alpha}}(\boldsymbol{x})=2C_{2}(n,d)(4K)^{\frac{d}{p}}\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}\tilde{\phi}_{\boldsymbol{\alpha}}\left(\sum_{j=1}^{d}\eta_{j}K^{j}\right)-C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}$
and obtain that
$\left|\phi_{\boldsymbol{\alpha}}\left(\frac{\boldsymbol{\eta}(p)}{K}\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\frac{\boldsymbol{\eta}(p)}{K}\right)\right|$ $=2C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}(4K)^{\frac{d}{p}}|\tilde{\phi}_{\boldsymbol{\alpha}}(p)-\xi_{\boldsymbol{\alpha},p}|$
(35) $\leq 4^{\frac{2d+p}{2p}}\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}C_{2}(n,d)N^{-2(n-1)}L^{-2(n-1)}.$
Then we obtain that
$\|\phi_{\boldsymbol{\alpha}}\left(\boldsymbol{\phi}_{2}(\boldsymbol{x})\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\boldsymbol{x}\right)\|_{W^{1,p}(\Omega_{\boldsymbol{m}_{*}})}=$ $\|\phi_{\boldsymbol{\alpha}}\left(\boldsymbol{\phi}_{2}(\boldsymbol{x})\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\boldsymbol{x}\right)\|_{L^{p}(\Omega_{\boldsymbol{m}_{*}})}$
(36) $\leq$ $4^{\frac{2d+p}{2p}}C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-1)}L^{-2(n-1)}$
which is due to $\phi_{\boldsymbol{\alpha}}\left(\boldsymbol{\phi}_{2}(\boldsymbol{x})\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\boldsymbol{x}\right)$ is a step function, and the first order weak derivative is 00 in $\Omega_{\boldsymbol{m}_{*}}$, and we have that
$\|\phi_{\boldsymbol{\alpha}}\left(\boldsymbol{\phi}_{2}(\boldsymbol{x})\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\boldsymbol{x}\right)\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}=$ $\|\phi_{\boldsymbol{\alpha}}\left(\boldsymbol{\phi}_{2}(\boldsymbol{x})\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\boldsymbol{x}\right)\|_{W^{1,p}(\Omega_{\boldsymbol{m}_{*}})}$
$=$ $\|\phi_{\boldsymbol{\alpha}}\left(\boldsymbol{\phi}_{2}(\boldsymbol{x})\right)-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}\left(\boldsymbol{x}\right)\|_{L^{p}(\Omega_{\boldsymbol{m}_{*}})}$
(37) $\leq$ $4^{\frac{2d+p}{2p}}C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-1)}L^{-2(n-1)}.$
Due to (v) in Lemma 4.3, there is a $\sigma_{2}$ neural network $\phi_{5,\boldsymbol{\alpha}}(\boldsymbol{x})$ with the width $N$ due to $N>n^{d}$ and depth $\left\lceil\log_{2}L\right\rceil$ due to $L>n$ such that
(38) $\phi_{5,\boldsymbol{\alpha}}(\boldsymbol{x})=\boldsymbol{x}^{\boldsymbol{\alpha}},~{}\boldsymbol{x}\in\mathbb{R}^{d}.$
Due to (iv) in Lemma 4.3, there is a $\sigma_{2}$ neural network $\phi_{6}(\boldsymbol{x})$ with the width $4$ and depth $1$ such that
(39) $\phi_{6}(x,y)=xy,~{}x,y\in\mathbb{R}.$
Now we define the neural network $\gamma_{\boldsymbol{m}_{*}}(\boldsymbol{x})$ to approximate $f_{K,{\boldsymbol{m}_{*}}}(\boldsymbol{x})$ in $\Omega_{\boldsymbol{m}_{*}}$:
(40) $\gamma_{\boldsymbol{m}_{*}}(\boldsymbol{x})=\sum_{|\boldsymbol{\alpha}|\leq n-1}\phi_{6}\left[\phi_{\boldsymbol{\alpha}}(\boldsymbol{\phi}_{2}(\boldsymbol{x})),\phi_{5,\boldsymbol{\alpha}}(\boldsymbol{x})\right].$
$\gamma_{\boldsymbol{m}_{*}}$ is a $\sigma_{2}$-NN with depth $\log_{2}9(L+1)$ and the width $28n^{d+1}(N+d)\log_{2}(8N)$.
The remaining question is to find the error $\mathcal{E}$:
$\widetilde{\mathcal{E}}:=$ $\left\|\sum_{|\boldsymbol{\alpha}|\leq n-1}\phi_{6}\left[\phi_{\boldsymbol{\alpha}}(\boldsymbol{\phi}_{2}(\boldsymbol{x})),\phi_{5,\boldsymbol{\alpha}}(\boldsymbol{x})\right]-f_{K,\boldsymbol{m}_{*}}(\boldsymbol{x})\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}$
$\leq$ $\sum_{|\boldsymbol{\alpha}|\leq n-1}\left\|\phi_{6}\left[\phi_{\boldsymbol{\alpha}}(\boldsymbol{\phi}_{2}(\boldsymbol{x})),\phi_{5,\boldsymbol{\alpha}}(\boldsymbol{x})\right]-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}$
$=$ $\sum_{|\boldsymbol{\alpha}|\leq n-1}\left\|\phi_{\boldsymbol{\alpha}}(\boldsymbol{\phi}_{2}(\boldsymbol{x}))\boldsymbol{x}^{\boldsymbol{\alpha}}-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}$
$\leq$ $n^{2}\sum_{|\boldsymbol{\alpha}|\leq n-1}\left\|\phi_{\boldsymbol{\alpha}}(\boldsymbol{\phi}_{2}(\boldsymbol{x}))-g_{f,\boldsymbol{\alpha},\boldsymbol{m}_{*}}(\boldsymbol{x})\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}$
(41) $\leq$ $n^{d+2}4^{\frac{2d+p}{2p}}C_{2}(n,d)\|f\|_{W^{n-1,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-1)}L^{-2(n-1)}.$
Combining Eqs. (33) and (41), we have that there is a DSRN $\gamma_{\boldsymbol{m}_{*}}\in\mathcal{N}_{2,9(L+1)\log_{2}L}$ with the width $28n^{d+1}(N+d)\log_{2}(8N)$ such that
$\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}_{*}}(\boldsymbol{x})\|_{W^{2,p}(\Omega_{\boldsymbol{m}_{*}})}\leq C_{10}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-2)/d}L^{-2(n-2)/d}$
$\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}_{*}}(\boldsymbol{x})\|_{W^{1,p}(\Omega_{\boldsymbol{m}_{*}})}\leq C_{10}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2(n-1)/d}L^{-2(n-1)/d}$
(42) $\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}_{*}}(\boldsymbol{x})\|_{L^{p}(\Omega_{\boldsymbol{m}_{*}})}\leq C_{10}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}_{*}})}N^{-2n/d}L^{-2n/d},$
where $C_{10}=C_{1}+4^{\frac{2d+p}{2p}}C_{2}(n,d)$ is the constant independent of $N,L$. The final inequality arises from the fact that $\frac{2n}{d}\leq 2(n-1)$ holds for values of $d$ and $n$ greater than or equal to 2.
Similarly, we can construct a $\gamma_{\boldsymbol{m}}$ with depth $\log_{2}9(L+1)$ and the width $28n^{d+1}(N+d)\log_{2}(8N)$ on $\Omega_{\boldsymbol{m}}$ with same order of Eq. (42).
According to Proposition 4.8, there is a sequence of the neural network $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ satisfies:
(i) $\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{L^{\infty}([a,b]^{d})}\leq 1$$\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{1,\infty}([a,b]^{d})}\leq\frac{8K}{b-a}$ and $\|s_{\boldsymbol{m}}(\boldsymbol{x})\|_{W^{2,\infty}([a,b]^{d})}\leq\frac{64K^{2}}{(b-a)^{2}}$ for any $\boldsymbol{m}\in\{1,2\}^{d}$.
(ii) $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$ is a partition of the unity $[a,b]^{d}$ with ${\rm supp}~{}s_{\boldsymbol{m}}(\boldsymbol{x})\cap[a,b]^{d}=\Omega_{\boldsymbol{m}}$ defined in Definition 4.5.
Due to (ii) in Lemma 4.3, there is a $\sigma_{2}$ neural network $\phi_{7}(x,y,z)$ with the width $96$ and depth $2$ such that
(43) $\phi_{7}(x,y,z)=xyz,~{}x,y,z\in\mathbb{R}.$
Now we define the piece-wise constant function $\phi_{\boldsymbol{\alpha}}(\boldsymbol{\phi}_{2}(\boldsymbol{x}))$ (37) defined on $\Omega_{\boldsymbol{m}}$ as $p_{\boldsymbol{m}}$, then we have
$\gamma(\boldsymbol{x})$ $=\sum_{\boldsymbol{m}\in\{1,2\}^{d}}s_{\boldsymbol{m}}(\boldsymbol{x})\cdot\gamma_{\boldsymbol{m}}(\boldsymbol{x})=\sum_{\boldsymbol{m}\in\{1,2\}^{d}}s_{\boldsymbol{m}}(\boldsymbol{x})\cdot\sum_{|\boldsymbol{\alpha}|\leq n-1}p_{\boldsymbol{m}}\cdot\boldsymbol{x}^{\boldsymbol{\alpha}}$
$=\sum_{|\boldsymbol{\alpha}|\leq n-1}\sum_{\boldsymbol{m}\in\{1,2\}^{d}}s_{\boldsymbol{m}}(\boldsymbol{x})\cdot p_{\boldsymbol{m}}\cdot\boldsymbol{x}^{\boldsymbol{\alpha}}$
$=\sum_{|\boldsymbol{\alpha}|\leq n-1}\sum_{\boldsymbol{m}\in\{1,2\}^{d}}(\lambda_{\boldsymbol{m},2}\circ\lambda_{\boldsymbol{m},1})(\boldsymbol{x})\cdot p_{\boldsymbol{m}}\cdot\boldsymbol{x}^{\boldsymbol{\alpha}}$
(44) $=\sum_{|\boldsymbol{\alpha}|\leq n-1}\sum_{\boldsymbol{m}\in\{1,2\}^{d}}\phi_{6}(\phi_{6}((\lambda_{\boldsymbol{m},2},\textbf{Id})\circ(\lambda_{\boldsymbol{m},1},p_{\boldsymbol{m}})),\boldsymbol{x}^{\boldsymbol{\alpha}}),$
where we denote $(f_{1},f_{2})\circ(g_{1},g_{2})$ as $(f_{1}\circ g_{1},f_{2}\circ g_{2})$. It is easy to check that $\gamma$ is a DSRN in $\mathcal{N}_{1,9(L+1)\log_{2}4L}$ based on Figure 4. Furthermore, the width of $\gamma_{\boldsymbol{m}_{*}}$ is $2^{d+6}n^{d+1}(N+d)\log_{2}(8N)$.
Figure 4: The structure of $\gamma(\boldsymbol{x})$.
Note that
$\widetilde{\mathcal{R}}:=$ $\|f(\boldsymbol{x})-\gamma(\boldsymbol{x})\|_{W^{2,p}([a,b]^{d})}\leq\sum_{\boldsymbol{m}\in\{1,2\}^{d}}\left\|s_{\boldsymbol{m}}(x)\cdot f(\boldsymbol{x})-s_{\boldsymbol{m}}(\boldsymbol{x})\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{W^{2,p}([a,b]^{d})}$
(45) $=$ $\sum_{\boldsymbol{m}\in\{1,2\}^{d}}\left\|s_{\boldsymbol{m}}(x)\cdot f(\boldsymbol{x})-s_{\boldsymbol{m}}(\boldsymbol{x})\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}})},$
where the last equality is due to ${\rm supp}~{}s_{\boldsymbol{m}}(\boldsymbol{x})\cap[a,b]^{d}=\Omega_{\boldsymbol{m}}$.
Then due to chain rule, for each $\boldsymbol{m}\in\{1,2\}^{d}$, we have
$\left\|s_{\boldsymbol{m}}(x)\cdot f(\boldsymbol{x})-s_{\boldsymbol{m}}(\boldsymbol{x})\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$
$\leq$ $\left\|s_{\boldsymbol{m}}(x)\right\|_{W^{2,\infty}(\Omega_{\boldsymbol{m}})}\left\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{L^{p}(\Omega_{\boldsymbol{m}})}+2\left\|s_{\boldsymbol{m}}(x)\right\|_{W^{1,\infty}(\Omega_{\boldsymbol{m}})}\left\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{W^{1,p}(\Omega_{\boldsymbol{m}})}$
$+\left\|s_{\boldsymbol{m}}(x)\right\|_{L^{\infty}(\Omega_{\boldsymbol{m}})}\left\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}+\left\|s_{\boldsymbol{m}}(x)\right\|_{W^{1,\infty}(\Omega_{\boldsymbol{m}})}\left\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{L^{p}(\Omega_{\boldsymbol{m}})}$
$+\left\|s_{\boldsymbol{m}}(x)\right\|_{L^{\infty}(\Omega_{\boldsymbol{m}})}\left\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{W^{1,p}(\Omega_{\boldsymbol{m}})}+\left\|s_{\boldsymbol{m}}(x)\right\|_{L^{\infty}(\Omega_{\boldsymbol{m}})}\left\|f(\boldsymbol{x})-\gamma_{\boldsymbol{m}}(\boldsymbol{x})\right\|_{L^{p}(\Omega_{\boldsymbol{m}})}$
(46) $\leq$ $91C_{10}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}N^{-2(n-2)/d}L^{-2(n-2)/d}\leq 91C_{10}(n,d)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d}.$
Hence
~$\widetilde{\mathcal{R}}\leq 2^{d+7}C_{10}(n,d)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-2)/d}L^{-2(n-2)/d}.$
Overall, we have shown that DSRNs can approximate functions in Sobolev spaces with the $W^{2,p}$ norm. Our method extends directly to $W^{m,p}$ for any integer $m\geq 1$. The key modification in the proof is the construction of a $C^{m}$ partition of unity $\{s_{\boldsymbol{m}}(\boldsymbol{x})\}_{\boldsymbol{m}\in\{1,2\}^{d}}$, which we achieve by designing each $s_{\boldsymbol{m}}(\boldsymbol{x})$ as a piecewise polynomial of sufficiently high degree. In particular, to build a $C^{3}$ partition of unity, one may take the one-dimensional “bump” function
[…]
This $C^{3}$ function admits an exact representation by $\mathrm{ReLU}^{5}$:
$s(x)\;=\;\sum_{k=0}^{5}(-1)^{k}\binom{5}{k}\,\Bigl{[}\tfrac{x}{3}+\tfrac{2}{3}-k\Bigr{]}_{+}^{5}.$
By Lemma 4.3, each $\mathrm{ReLU}^{5}$ can itself be realized by a $\sigma_{2}$-NN. Consequently, we obtain a DSRN that approximates the target function in the $W^{3,p}$ norm. Repeating this construction inductively yields Corollary 4.23.
Corollary 4.23.
For any $f\in W^{n,p}([a,b]^{d})$, $m\in\mathbb{N}$ with $m\geq 2$ and $1\leq p\leq+\infty$, any $N,L\in\mathbb{N}_{+}$ with $N\log_{2}L+2^{\left\lfloor\log_{2}N\right\rfloor}\geq\max\{d,n\}$ and $L\geq N,m<n$, there is a DSRN $\gamma(\boldsymbol{x})$ in $\mathcal{N}_{\eta_{1},\eta_{2}L\log_{2}L}$ with the width $\eta_{3}N\log_{2}N$ such that
$\|f(\boldsymbol{x})-\gamma(\boldsymbol{x})\|_{W^{m,p}([a,b]^{d})}\leq C_{11}(n,d)\|f\|_{W^{n,p}([a,b]^{d})}N^{-2(n-m)/d}L^{-2(n-m)/d},$
where $\eta_{i},C_{11}$ are the constants independent of $N,L$.
In the above proof, we remove the non-smooth points using a smooth partition of unity. In numerical practice, even if a DSRN with a random initial condition does not contain second-order derivatives across the entire domain, it does have second-order derivatives at most points, meaning that the non-smooth points are of zero measure. If we unfortunately sample points at the non-smooth locations, we can resample in the next iteration to avoid these points. In Sec. 7, we demonstrate that DSRN works effectively in practice.
5 Optimality of Theorem 4.1
In this section, our aim is to demonstrate the optimality of DSRN approximations in Sobolev spaces measured by norms in $W^{m,p}\left([a,b]^{d}\right)$ for $m\geq 2$. Without loss of generality, we prove the optimality for $m=2$.
Theorem 5.1.
Given any $\rho,C_{1},C_{2},C_{3},C_{4},J_{0}>0$ and $n,d\in\mathbb{N}^{+}$, there exist $N,L\in\mathbb{N}$ with $NL\geq J_{0}$ and $f$, satisfying for any $\phi\in\mathcal{N}_{C_{4},C_{2}L\log L}$ with the width smaller than $C_{1}N\log N$, we have
(47) $|\phi-f|_{W^{2,p}([a,b]^{d})}>C_{3}\|f\|_{W^{n,p}([a,b]^{d})}L^{-2(n-2)/d-\rho}N^{-2(n-2)/d-\rho}.$
In order to prove the optimality of Theorem 4.1, we need to bound the VC-dimension of second order derivative of deep neural networks (DNNs).
Definition 5.2 (VC-dimension [1]).
Let $H$ denote a class of functions from $\mathcal{X}$ to $\{0,1\}$. For any non-negative integer $m$, define the growth function of $H$ as
$\Pi_{H}(m):=\max_{x_{1},x_{2},\ldots,x_{m}\in\mathcal{X}}\left|\{\left(h(x_{1}),h(x_{2}),\ldots,h(x_{m})\right):h\in H\}\right|.$
The Vapnik–Chervonenkis dimension (VC-dimension) of $H$, denoted by $\text{VCdim}(H)$, is the largest $m$ such that $\Pi_{H}(m)=2^{m}$. For a class $\mathcal{G}$ of real-valued functions, define $\text{VCdim}(\mathcal{G}):=\text{VCdim}(\mathrm{sgn}(\mathcal{G}))$, where $\mathrm{sgn}(\mathcal{G}):=\{\mathrm{sgn}(f):f\in\mathcal{G}\}$ and $\mathrm{sgn}(x)=1[x>0]$.
Theorem 5.3.
For any $N,L,d,C\in\mathbb{N}_{+}$, there exists a constant $\bar{C}$ independent of $N,L$ such that
(48) $\text{VCdim}(D^{2}\Phi)\leq\bar{C}N^{2}L^{2}\log_{2}L\log_{2}N,$
for
(49) $D^{2}\Phi:=\left\{\lambda(\boldsymbol{x})=D^{\boldsymbol{\alpha}}\phi:\phi\in\Phi,~{}|\boldsymbol{\alpha}|=2\right\},$
where $\Phi:=\left\{\phi:\phi\text{ is a DSRN in $\mathcal{N}_{C,L}$ with width$\leq N$},\phi\in W^{2}([a,b]^{d})\right\}$, and
$D^{\boldsymbol{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$
as the partial derivative where $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the weak derivative in the $i$-th variable.
Prior to proving Theorem 5.3, we introduce two lemmas.
Lemma 5.4 ([3],[2]).
Suppose $W\leq M$ and let $P_{1},\ldots,P_{M}$ be polynomials of degree at most $D$ in $W$ variables. Define
$P:=\left|\{\left(\mathrm{sgn}(P_{1}(\boldsymbol{a})),\ldots,\mathrm{sgn}(P_{M}(\boldsymbol{a}))\right):\boldsymbol{a}\in\mathbb{R}^{W}\}\right|,$
then we have $P\leq 2(2eMD/W)^{W}$.
Lemma 5.5 ([3]).
Suppose that $2^{m}\leq 2^{t}(mr/w)^{w}$ for some $r\geq 16$ and $m\geq w\geq t\geq 0$. Then, $m\leq t+w\log_{2}(2r\log_{2}r)$.
Proof 5.6 (Proof of Theorem 5.1).
The difficulty of the estimate raises from the complexity of the second order derivatives of DNNs:
$\phi=\boldsymbol{W}_{L+1}\sigma(\boldsymbol{W}_{L}\sigma(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})+b_{L+1},$
where $\sigma$ can be either the ReLU or the ReLU square. Then the first order derivative can be read as
$\psi(\boldsymbol{x})=D_{i}\phi(\boldsymbol{x})=$ $\boldsymbol{W}_{L+1}\sigma^{\prime}(\boldsymbol{W}_{L}\sigma(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})$
(50) $\cdot\boldsymbol{W}_{L}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\ldots\boldsymbol{W}_{2}\sigma^{\prime}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})(\boldsymbol{W}_{1})_{i},$
where $\boldsymbol{W}_{i}\in\mathbb{R}^{N_{i}\times N_{i-1}}$ ($(\boldsymbol{W})_{i}$ is $i$-th column of $\boldsymbol{W}$) and $\boldsymbol{b}_{i}\in\mathbb{R}^{N_{i}}$ are the weight matrix and the bias vector in the $i$-th linear transform in $\phi$, and $\sigma^{\prime}(\boldsymbol{x})=\operatorname{diag}(\sigma^{\prime}(x_{i}))$. Then we have
(51) $\lambda(\boldsymbol{x})=D_{j}D_{i}\phi(\boldsymbol{x})=\sum_{a=1}^{L+1}\lambda_{a}(\boldsymbol{x}).$
For
$\lambda_{a}(\boldsymbol{x})=$ $\boldsymbol{W}_{L+1}\sigma^{\prime}(\boldsymbol{W}_{L}\sigma(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})\cdot\boldsymbol{W}_{L}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\ldots$
$\cdot[\boldsymbol{W}_{a}\sigma^{\prime\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\cdot\boldsymbol{W}_{a-1}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\boldsymbol{W}_{2}\sigma^{\prime}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})(\boldsymbol{W}_{1})_{j}]$
(52) $\cdot\ldots\boldsymbol{W}_{a-1}\sigma^{\prime}(\ldots\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)\boldsymbol{W}_{2}\sigma^{\prime}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})(\boldsymbol{W}_{1})_{i},$
where $\sigma^{\prime\prime}(\boldsymbol{x})=\operatorname{diag}(\sigma^{\prime\prime}(x_{i}))$ is a three-order tensor. Denote $W_{i}$ as the number of parameters in $\boldsymbol{W}_{i},\boldsymbol{b}_{i}$, i.e., $W_{i}=N_{i}N_{i-1}+N_{i}$.
Figure 5: The structure of $\lambda_{a}(\boldsymbol{x})$.
Note that $\sigma$ can either be the ReLU or the ReLU squared function. If $\sigma(x)$ is the ReLU function, then its second derivative, $\sigma^{\prime\prime}(x)$, will be the delta function, which is defined as $\sigma^{\prime\prime}(x)=0$ for $x\neq 0$ and $\sigma^{\prime\prime}(0)=\infty$, with $\int_{\mathbb{R}}\sigma^{\prime\prime}(x)\,\mathrm{d}x=1$. This delta function serves as the generalization function. However, since we assume that $\phi\in W^{2}([a,b]^{d})$, any delta functions must either disappear or be omitted in the $\lambda_{a}(\boldsymbol{x})$ equation, as shown in Eq. (52).
Let $\boldsymbol{x}\in\mathbb{R}^{d}$ be an input and $\boldsymbol{\theta}\in\mathbb{R}^{W}$ be a parameter vector in $\psi$. We denote the output of $\psi$ with input $\boldsymbol{x}$ and parameter vector $\boldsymbol{\theta}$ as $f(\boldsymbol{x},\boldsymbol{\theta})$. For fixed $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{m}$ in $\mathbb{R}^{d}$, we aim to bound
(53) $K:=\left|\{\left(\mathrm{sgn}(f(\boldsymbol{x}_{1},\boldsymbol{\theta})),\ldots,\mathrm{sgn}(f(\boldsymbol{x}_{m},\boldsymbol{\theta}))\right):\boldsymbol{\theta}\in\mathbb{R}^{W}\}\right|.$
The proof is follow our previous work [42]. For any partition $\mathcal{S}=\{P_{1},P_{2},\ldots,P_{T}\}$ of the parameter domain $\mathbb{R}^{W}$, we have
$K\leq\sum_{i=1}^{T}\left|\{\left(\mathrm{sgn}(f(\boldsymbol{x}_{1},\boldsymbol{\theta})),\ldots,\mathrm{sgn}(f(\boldsymbol{x}_{m},\boldsymbol{\theta}))\right):\boldsymbol{\theta}\in P_{i}\}\right|.$
We choose the partition such that within each region $P_{i}$, the functions $f(\boldsymbol{x}_{j},\cdot)$ are all fixed polynomials of bounded degree. This allows us to bound each term in the sum using Lemma 5.4.
We define a sequence of sets of functions $\{\mathbb{F}_{j}\}_{j=0}^{L}$ with respect to parameters $\boldsymbol{\theta}\in\mathbb{R}^{W}$:
[…]
$\boldsymbol{W}_{3}\sigma(\boldsymbol{W}_{2}\sigma(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})+\boldsymbol{b}_{2})+\boldsymbol{b}_{3}\}\cup\mathbb{F}_{1}$
$\vdots$
[…]
(54) $\boldsymbol{W}_{L+1}\sigma^{\prime\prime}(\boldsymbol{W}_{L}\sigma_{1}(\ldots\sigma_{1}(\boldsymbol{W}_{1}\boldsymbol{x}+\boldsymbol{b}_{1})\ldots)+\boldsymbol{b}_{L})\}\cup\mathbb{F}_{L-1}.$
The partition of $\mathbb{R}^{W}$ is constructed layer by layer through successive refinements denoted by $\mathcal{S}_{0},\mathcal{S}_{1},\ldots,\mathcal{S}_{L}$. We denote $L^{*}=L-C\log_{2}L$. These refinements possess the following properties:
1. We have $|\mathcal{S}_{0}|=1$, and for each $n=1,\ldots,L$, we have
$\frac{|\mathcal{S}_{n}|}{|\mathcal{S}_{n-1}|}\leq 2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}.$
2. For each $n=0,\ldots,L^{*}$, each element $S$ of $\mathcal{S}_{n}$, when $\boldsymbol{\theta}$ varies in $S$, the output of each term in $\mathbb{F}_{n}$ is a fixed polynomial function in $\sum_{i=1}^{n}W_{i}$ variables of $\boldsymbol{\theta}$, with a total degree no more than $1+n2^{\max\{0,n-L^{*}\}}$.
We define $\mathcal{S}_{0}=\{\mathbb{R}^{W}\}$, which satisfies properties 1,2 above, since $\boldsymbol{W}_{1}\boldsymbol{x}_{j}+\boldsymbol{b}_{1}$ and $(\boldsymbol{W}_{1})_{i}$ for all $i=1,\ldots,d$ are affine functions of $\boldsymbol{W}_{1},\boldsymbol{b}_{1}$.
For each $n=0,\ldots,L$, to define $\mathcal{S}_{n}$, we use the last term of $\mathbb{F}_{n-1}$ as inputs for the new terms in $\mathbb{F}_{n}$. Assuming that $\mathcal{S}_{0},\mathcal{S}_{1},\ldots,\mathcal{S}_{n-1}$ have already been defined, we observe that the last two or three terms are new additions to $\mathbb{F}_{n}$ when comparing it to $\mathbb{F}_{n-1}$. Therefore, all elements in $\mathbb{F}_{n}$ except the $\mathbb{F}_{n}\backslash\mathbb{F}_{n-1}$ are fixed polynomial functions in $W_{n}$ variables of $\boldsymbol{\theta}$, with a total degree no greater than $1+(n-1)2^{\max\{0,n-1-L^{*}\}}$ when $\boldsymbol{\theta}$ varies in $S\in\mathcal{S}_{n}$. This is because $\mathcal{S}_{n}$ is a finer partition than $\mathcal{S}_{n-1}$.
We denote $p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta})$ as the output of the $k$-th node in the last term of $\mathbb{F}_{n-1}$ in response to $\boldsymbol{x}_{j}$ when $\boldsymbol{\theta}\in S$. The collection of polynomials
$\{p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta}):j=1,\ldots,m,~{}k=1,\ldots,N_{n}\}$
can attain at most $2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}$ distinct sign patterns when $\boldsymbol{\theta}\in S$ due to Lemma 5.4 for sufficiently large $m$. Therefore, we can divide $S$ into
$2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}$
parts, each having the property that $p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta})$ does not change sign within the subregion. By performing this for all $S\in\mathcal{S}_{n-1}$, we obtain the desired partition $\mathcal{S}_{n}$. This division ensures that the required property 1 is satisfied.
Additionally, since the input to the last two terms in $\mathbb{F}_{n}$ is $p_{\boldsymbol{x}_{j},n-1,S,k}(\boldsymbol{\theta})$, and we have shown that the sign of this input will not change in each region of $\mathcal{S}_{n}$, it follows that the output of the last two terms in $\mathbb{F}_{n}$ is also a polynomial without breakpoints in each element of $\mathcal{S}_{n}$, therefore, the required property 2 is satisfied.
Due to the structure of $\lambda$, it is a polynomial function in $\sum_{i=1}^{L+1}W_{i}$ variables of $\boldsymbol{\theta}\in S\in\mathcal{S}_{L}$, of total degree no more than
$d_{2}:=2\sum_{n=0}^{L}(1+n2^{\max\{0,n-L^{*}\}})=2L+2+L^{C}(L-1).$
Therefore, for each $S\in\mathcal{S}_{L}$ we have
|$\left|\{\left(\mathrm{sgn}(f(\boldsymbol{x}_{1},\boldsymbol{\theta})),\ldots,\mathrm{sgn}(f(\boldsymbol{x}_{m},\boldsymbol{\theta}))\right):\boldsymbol{\theta}\in S\}\right|\leq 2\left(2emd_{2}/\sum_{i=1}^{L+1}W_{i}\right)^{\sum_{i=1}^{L+1}W_{i}}.$
Then
$K\leq$ $2\left(2emd_{2}/\sum_{i=1}^{L+1}W_{i}\right)^{\sum_{i=1}^{L+1}W_{i}}\cdot\prod_{n=1}^{L}2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}$
$\leq$ $4\prod_{n=1}^{L+1}2\left(\frac{2em(1+(n-1)2^{\max\{0,n-1-L^{*}\}})N_{n}}{\sum_{i=1}^{n}W_{i}}\right)^{\sum_{i=1}^{n}W_{i}}$
(55) $\leq$ $2^{L+3}\left(\frac{4em(L^{C+3})N}{U}\right)^{U},$
where $U:=\sum_{n=1}^{L+1}\sum_{i=1}^{n}W_{i}=\mathcal{O}(N^{2}L^{2})$, $N$ is the width of the network, and the last inequality is due to weighted AM-GM. For the definition of the VC-dimension, we have
(56) $2^{\text{VCdim}(D^{2}\Phi)}\leq 2^{L+3}\left(\frac{4e\text{VCdim}(D^{2}\Phi)L^{C+3}N}{U}\right)^{U}.$
Due to Lemma 5.5, we obtain that
(57) $\text{VCdim}(D^{2}\Phi)=\mathcal{O}(N^{2}L^{2}\log_{2}L\log_{2}N)$
since $U=\mathcal{O}(N^{2}L^{2})$.
Our method for estimating the VC-dimension is also suitable for the higher-order derivatives of DNNs. The reason for this is that the partition of $\mathcal{S}$ and the definition of the sets of function sets $\{\mathbb{F}_{j}\}_{j=0}^{L}$ do not depend on the order of the derivatives in our method. Therefore, the VC-dimension of the higher-order derivatives of DSRNs is still $\mathcal{O}(N^{2}L^{2}\log_{2}L\log_{2}N)$.
Remark 5.7.
The estimation of the VC-dimension of higher-order derivatives of deep neural networks can follow our proof above. Based on this proof, we can notice that if the square of ReLU can appear at any position instead of just the last several layers like in DSRN, the upper bound of the VC-dimension will be $\mathcal{O}(N^{2}L^{3}\log N\log L)$, which is much more complex than DSRN. This will make the proof of optimality to invalid. Furthermore, this can be one reason to show that deep neural networks with only the square of ReLU as the activation function are more difficult to train since a larger VC-dimension means increased complexity of the space.
Proof 5.8 (Proof of Theorem 5.1).
By combining Theorem 5.3 with the result presented in [34, Theorem 5], we can establish the proof for Theorem 5.1.
Furthermore, through the combination of Theorem 4.1 with the findings outlined in [34, Theorem 5], we can establish the following corollary. This corollary serves to demonstrate the optimality stated in Theorem 5.3.
Corollary 5.9.
For any $d\in\mathbb{N}_{+}$, $C,J_{0},\varepsilon>0$, there exists $N,L\in\mathbb{N}$ with $NL\geq J_{0}$ such that
(58) $\text{VCdim}(D^{2}\Phi)>CN^{2-\varepsilon}L^{2-\varepsilon},$
where $D^{2}\Phi$ is defined in Theorem 5.3.
Note that in this section, we have proven the optimality of the approximation of DSRNs. However, if we allow the square of ReLU to appear in any layer of DNNs, the VC-dimension of derivatives of such DNNs will become $\mathcal{O}(L^{3}N^{2})$, as also shown in [3]. In other words, we cannot prove the optimality of $\sigma_{2}$-NNs based on such a bound of VC-dimension. Nevertheless, we believe that the approximation rate $\mathcal{O}(N^{-2(n-2)/d}L^{-2(n-2)/d})$ may not be improved to $\mathcal{O}(N^{-2(n-2)/d}L^{-3(n-2)/d})$ for $\sigma_{2}$-NNs, as ReLU can efficiently approximate the square of ReLU [43], and the optimality of ReLU DNNs has been proven in [26, 42]. Obtaining optimal bounds for $\sigma_{2}$-NNs will be a future research direction.
6 Generalization Analysis in Sobolev Spaces
In a typical supervised learning algorithm, the objective is to learn a high-dimensional target function $f(\boldsymbol{x})$ defined on $[a,b]^{d}$ with $\|f\|_{W^{n,\infty}\left([a,b]^{d}\right)}\leq B$ for $B>0$ from a finite set of data samples $\{(\boldsymbol{x}_{i},f(\boldsymbol{x}_{i}))\}_{i=1}^{M}$. When training a DNN, we aim to identify a DNN $\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})$ that approximates $f(\boldsymbol{x})$ based on random data samples $\{(\boldsymbol{x}_{i},f(\boldsymbol{x}_{i}))\}_{i=1}^{M}$. We assume that $\{\boldsymbol{x}_{i}\}_{i=1}^{M}$ is an i.i.d. sequence of random variables uniformly distributed on $[a,b]^{d}$ in this section. Denote
(59) $\boldsymbol{\theta}_{D}$ $:=\arg\inf_{\boldsymbol{\theta}}\mathcal{R}_{D}(\boldsymbol{\theta}):=\arg\inf_{\boldsymbol{\theta}}\int_{[a,b]^{d}}|\Delta(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}))|^{2}+|f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta})|^{2}\,\mathrm{d}\boldsymbol{x},$
(60) $\boldsymbol{\theta}_{S}$ $:=\arg\inf_{\boldsymbol{\theta}}\mathcal{R}_{S}(\boldsymbol{\theta}):=\arg\inf_{\boldsymbol{\theta}}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left[|\Delta(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}))|^{2}+|f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta})|^{2}\right].$
The overall inference error is $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})$, which can be divided into two parts:
$\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})=$ $\mathcal{R}_{D}(\boldsymbol{\theta}_{D})+\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{D})-\mathcal{R}_{D}(\boldsymbol{\theta}_{D})+\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{D})+\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})$
(61) $\leq$ $\underbrace{\mathcal{R}_{D}(\boldsymbol{\theta}_{D})}_{\text{approximation error}}+\underbrace{\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S}),}_{\text{{generalization error}}}$
where the last inequality is due to $\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\leq\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{D})$ by the definition of $\boldsymbol{\theta}_{S}$ and the definition of the integral.
Due to Theorem 4.1, we know that the approximation error
$\mathcal{R}_{D}(\boldsymbol{\theta}_{D})\sim\mathcal{O}(N^{-4(n-2)/d}L^{-4(n-2)/d})$
term since $\|f(\boldsymbol{x})-\phi(\boldsymbol{x})\|_{H^{2}([a,b]^{d})}\leq\|f(\boldsymbol{x})-\phi(\boldsymbol{x})\|_{W^{2,\infty}([a,b]^{d})}$. In this section, we bound generalization error in the $H^{1}([a,b]^{d})$ sense. We will bound this error in two ways. The first approach is based on the Rademacher complexity, which provides an error bound that is more favorable with respect to the number of parameters (Theorem 6.1). The second approach utilizes the inequality from [16, Theorem 11.4], which is more favorable in terms of the number of sample points $M$ (Theorem 6.2).
Theorem 6.1.
For any $N,L,d,B,C_{1},C_{2},C_{3}$, if $\phi(\boldsymbol{x};\boldsymbol{\theta}_{D}),\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\in\widetilde{\Phi}$, we will have that there are constants $C_{5}=C_{5}(B,d,C_{1},C_{2},C_{3})$ and $J=J(d,N,L,C_{1},C_{2},C_{3})$ such that for any $M\geq J$, we have
(62) $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})\leq C_{5}\left[\left(NL\right)^{-\frac{4(n-2)}{d}}+\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M\right]$
where $\widetilde{\Phi}:=\{\phi:\text{$\phi\in\mathcal{N}_{C_{1},C_{2}L\log_{2}L}$ with the width $\leq C_{3}N\log N$, }\|\phi\|_{W^{2,\infty}([a,b]^{d})}\leq B\}$, and $\mathcal{R}_{S},\mathcal{R}_{D},\boldsymbol{\theta}_{S},\boldsymbol{\theta}_{D}$ are defined in Eqs. (59,60).
The proof of Theorem 6.1 is based on the works of [2, 7, 24, 42]. We begin by bounding the generalization error using the Rademacher Complexity and then bound the Rademacher Complexity by the uniform covering number. We further bound the uniform covering number by the pseudo-dimension. Finally, we estimate the pseudo-dimension by Theorem 6.13.
Theorem 6.2.
For any $N,L,d,B,C_{1},C_{2},C_{3}$, if $\phi(\boldsymbol{x};\boldsymbol{\theta}_{D}),\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\in\widetilde{\Phi}$, we will have that there are constants $C_{6}=C_{6}(B,d,C_{1},C_{2},C_{3})$ and $J=J(d,N,L,C_{1},C_{2},C_{3})$ such that for any $M\geq J$, we have
(63) $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})\leq C_{6}\left[\left(NL\right)^{-\frac{4(n-2)}{d}}+\frac{N^{2}L^{2}(\log_{2}L\log_{2}N)}{M}\log M\right]$
where $\widetilde{\Phi}:=\{\phi:\text{$\phi\in\mathcal{N}_{C_{1},C_{2}L\log_{2}L}$ with the width $\leq C_{3}N\log N$, }\|\phi\|_{W^{2,\infty}([a,b]^{d})}\leq B\}$, and $\mathcal{R}_{S},\mathcal{R}_{D},\boldsymbol{\theta}_{S},\boldsymbol{\theta}_{D}$ are defined in Eqs. (59,60).
The proof of Theorem 6.2 relies on the inequality from [16, Theorem 11.4] and the estimation of the pseudo-dimension provided in Theorem 6.13.
Remark 6.3.
All the outcomes within this section can be naturally extended to scenarios involving measurements by norms that incorporate higher-order derivatives. The order of the interior error is given by:
$\mathcal{O}\left(\left(NL\right)^{-\frac{4(n-k)}{d}}+\frac{N^{2}L^{2}(\log_{2}L\log_{2}N)}{M}\log M\right)$
or
$\mathcal{O}\left(\left(NL\right)^{-\frac{4(n-k)}{d}}+\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M\right)$
where $k$ is the regularity of the loss functions.
6.1 Proof of Theorem 6.1
6.1.1 Bounding generalization error by Rademacher complexity
Definition 6.4 (Rademacher complexity [2]).
Given a sample set $S=\{z_{1},z_{2},\ldots,z_{M}\}$ on a domain $\mathcal{Z}$, and a class $\mathcal{F}$ of real-valued functions defined on $\mathcal{Z}$, the empirical Rademacher complexity of $\mathcal{F}$ in $S$ is defined as
$\mathbf{R}_{S}(\mathcal{F}):=\frac{1}{M}\mathbb{E}_{\Sigma_{M}}\left[\sup_{f\in\mathcal{F}}\sum_{i=1}^{M}\sigma_{i}f(z_{i})\right],$
where $\Sigma_{M}:=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{M}\}$ are independent random variables drawn from the Rademacher distribution, i.e., $\mathbb{P}(\sigma_{i}=+1)=\mathbb{P}(\sigma_{i}=-1)=\frac{1}{2}$ for $i=1,2,\ldots,M.$ For simplicity, if $S=\{z_{1},z_{2},\ldots,z_{M}\}$ is an independent random variable set with the uniform distribution, denote
$\mathbf{R}_{M}(\mathcal{F}):=\mathbb{E}_{S}\mathbf{R}_{S}(\mathcal{F}).$
The following lemma will be used to bounded generalization error by Rademacher complexities:
Lemma 6.5 ([38], Proposition 4.11).
Let $\mathcal{F}$ be a set of functions. Then
$\mathbb{E}_{X}\sup_{u\in\mathcal{F}}\left|\frac{1}{M}\sum_{i=1}^{M}u(x_{j})-\mathbb{E}_{x\sim\mathcal{P}_{\Omega}}u(x)\right|\leq 2\mathbf{R}_{M}(\mathcal{F}),$
where $X:=\{x_{1},\ldots,x_{M}\}$ is an independent random variable set with the uniform distribution.
Now we can show that generalization error can be bounded by Rademacher complexities of two function sets.
Lemma 6.6.
Let $d,N,L,M\in\mathbb{N}_{+}$, $B,C_{1},C_{2},C_{3}\in\mathbb{R}_{+}$. For any $f\in W^{2,\infty}([a,b]^{d})$ with $\|f\|_{W^{2,\infty}([a,b]^{d})}\leq B$, set
$\widetilde{\Phi}$ :=$:=\{\phi:\text{$\phi\in\mathcal{N}_{C_{1},C_{2}L\log_{2}L}$ with the width $\leq C_{3}N\log N$, }\|\phi\|_{W^{2,\infty}([a,b]^{d})}\leq B\}$
(64) $D^{2}\widetilde{\Phi}$ :=$:=\{\psi:\psi=D_{j}D_{i}\phi,i,j=1,\ldots,d\}.$
If $\phi(\boldsymbol{x};\boldsymbol{\theta}_{D}),\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\in\widetilde{\Phi}$, we have
$\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\leq 4B(d\mathbf{R}_{M}(D^{2}\widetilde{\Phi})+\mathbf{R}_{M}(\widetilde{\Phi})),$
where $\mathbb{E}$ is expected responding to $X$, and $X:=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{M}\}$ is an independent random variables set uniformly distributed on $[a,b]^{d}$.
Proof 6.7.
$\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})-\mathcal{R}_{D}(\boldsymbol{\theta}_{S})$
$=$ $\sum_{j=1}^{d}\left(\mathbb{E}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\frac{\partial^{2}(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}\right|^{2}-\int_{[a,b]^{d}}\left|\frac{\partial^{2}(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}\right|^{2}\,\mathrm{d}\boldsymbol{x}\right)$
$+\mathbb{E}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))\right|^{2}-\int_{[a,b]^{d}}\left|(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))\right|^{2}\,\mathrm{d}\boldsymbol{x}$
$\leq$ $2B\sum_{j=1}^{d}\left(\mathbb{E}\left|\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\frac{\partial^{2}(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}-\int_{[a,b]^{d}}\frac{\partial^{2}(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))}{\partial x_{j}^{2}}\,\mathrm{d}\boldsymbol{x}\right|\right)$
$+2B\mathbb{E}\left|\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}(f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S}))-\int_{[a,b]^{d}}(f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S}))\,\mathrm{d}\boldsymbol{x}\right|$
(65) $\leq$ $4B(d\mathbf{R}_{M}(D^{2}\widetilde{\Phi})+\mathbf{R}_{M}(\widetilde{\Phi}))$
where the last inequality is due to Lemma 6.6.
6.1.2 Bounding the Rademacher complexity and the proof of Theorem 6.1
In this subsection, we aim to estimate the Rademacher complexity using the covering number. We then estimate the covering number using the pseudo-dimension.
Definition 6.8 (covering number [2]).
Let $(V,\|\cdot\|)$ be a normed space, and $\Theta\in V$. $\{V_{1},V_{2},\ldots,V_{n}\}$ is an $\varepsilon$-covering of $\Theta$ if $\Theta\subset\cup_{i=1}^{n}B_{\varepsilon,\|\cdot\|}(V_{i})$. The covering number $\mathcal{N}(\varepsilon,\Theta,\|\cdot\|)$ is defined as
$\mathcal{N}(\varepsilon,\Theta,\|\cdot\|):=\min\{n:\exists\varepsilon\text{-covering over }\Theta\text{ of size }n\}\text{. }$
Definition 6.9 (Uniform covering number [2]).
Suppose the $\mathcal{F}$ is a class of functions from $\mathcal{F}$ to $\mathbb{R}$. Given $n$ samples $\boldsymbol{Z}_{n}=(z_{1},\ldots,z_{n})\in\mathcal{X}^{n}$, define
$\mathcal{F}|_{\boldsymbol{Z}_{n}}=\{(u(z_{1}),\ldots,u(z_{n})):u\in\mathcal{F}\}.$
The uniform covering number $\mathcal{N}(\varepsilon,\mathcal{F},n)$ is defined as
$\mathcal{N}(\varepsilon,\mathcal{F},n)=\max_{\boldsymbol{Z}_{n}\in\mathcal{X}^{n}}\mathcal{N}\left(\varepsilon,\mathcal{F}|_{\boldsymbol{Z}_{n}},\|\cdot\|_{\infty}\right),$
where $\mathcal{N}\left(\varepsilon,\mathcal{F}|_{\boldsymbol{Z}_{n}},\|\cdot\|_{\infty}\right)$ denotes the $\varepsilon$-covering number of $\mathcal{F}|_{\boldsymbol{Z}_{n}}$ w.r.t the $L_{\infty}$-norm.
Then we use a lemma to estimate the Rademacher complexity using the covering number.
Lemma 6.10 (Dudley’s theorem [2]).
Let $\mathcal{F}$ be a function class such that $\sup_{f\in\mathcal{F}}\|f\|_{\infty}\leq B$. Then the Rademacher complexity $\mathbf{R}_{n}(\mathcal{F})$ satisfies that
$\mathbf{R}_{n}(\mathcal{F})\leq\inf_{0\leq\delta\leq B}\left\{4\delta+\frac{12}{\sqrt{n}}\int_{\delta}^{B}\sqrt{\log 2\mathcal{N}(\varepsilon,\mathcal{F},n)}\,\mathrm{d}\varepsilon\right\}$
To bound the Rademacher complexity, we employ Lemma 6.10, which bounds it by the uniform covering number. We estimate the uniform covering number by the pseudo-dimension based on the following lemma.
Definition 6.11 (pseudo-dimension [30]).
Let $\mathcal{F}$ be a class of functions from $\mathcal{X}$ to $\mathbb{R}$. The pseudo-dimension of $\mathcal{F}$, denoted by $\text{Pdim}(\mathcal{F})$, is the largest integer $m$ for which there exists $(x_{1},x_{2},\ldots,x_{m},y_{1},y_{2},\ldots,y_{m})\in\mathcal{X}^{m}\times\mathbb{R}^{m}$ such that for any $(b_{1},\ldots,b_{m})\in\{0,1\}^{m}$ there is $f\in\mathcal{F}$ such that $\forall i:f\left(x_{i}\right)>y_{i}\Longleftrightarrow b_{i}=1.$
Lemma 6.12 ([2]).
Let $\mathcal{F}$ be a class of functions from $\mathcal{X}$ to $[-B,B]$. For any $\varepsilon>0$, we have
$\mathcal{N}(\varepsilon,\mathcal{F},n)\leq\left(\frac{2enB}{\varepsilon\text{Pdim}(\mathcal{F})}\right)^{\text{Pdim}(\mathcal{F})}$
for $n\geq\text{Pdim}(\mathcal{F})$.
The remaining problem is to bound $\text{Pdim}(\widetilde{\Phi})$ and $\text{Pdim}(D^{2}\widetilde{\Phi})$. Based on [3], $\text{Pdim}(\widetilde{\Phi})=O(L^{2}N^{2}\log_{2}L\log_{2}N)$. For the $\text{Pdim}(D^{2}\widetilde{\Phi})$, we can estimate it by Theorem 6.13.
Theorem 6.13.
For any $N,L,d,C\in\mathbb{N}_{+}$, there exists a constant $\bar{C}$ independent of $N,L$ such that
(66) $\text{Pdim}(D^{2}\Phi)\leq\bar{C}N^{2}L^{2}\log_{2}L\log_{2}N,$
for
(67) $D^{2}\Phi:=\left\{\lambda(\boldsymbol{x})=D^{\boldsymbol{\alpha}}\phi:\phi\in\Phi,~{}|\boldsymbol{\alpha}|=2\right\},$
where $\Phi:=\left\{\phi:\phi\text{ is a DSRN in $\mathcal{N}_{C,L}$ with width$\leq N$},\phi\in W^{2}([a,b]^{d})\right\}$, and
$D^{\boldsymbol{\alpha}}=D^{\alpha_{1}}_{1}D^{\alpha_{2}}_{2}\ldots D^{\alpha_{d}}_{d}$
as the partial derivative where $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{d}]^{T}$ and $D_{i}$ is the weak derivative in the $i$-th variable.
Proof 6.14.
Denote $D^{2}\Phi_{\mathcal{N}}:=\{\eta(\boldsymbol{x},y):\eta(\boldsymbol{x},y)=\psi(\boldsymbol{x})-y,\psi\in D^{2}\Phi,(\boldsymbol{x},y)\in\mathbb{R}^{d+1}\}.$ Based on the definition of VC-dimension and pseudo-dimension, we have that
(68) $\text{Pdim}(D^{2}\Phi)\leq\text{VCdim}(D^{2}{\Phi}_{\mathcal{N}}).$
For the $\text{VCdim}(D^{2}{\Phi}_{\mathcal{N}})$, it can be bounded by $O(N^{2}L^{2}\log_{2}L\log_{2}N)$. The proof is similar to that for the estimate of $\text{VCdim}(D^{2}{\Phi})$ as given in Theorem 5.3.
Now we can estimate generalization error based on Lemma 6.6.
Proof 6.15 (Proof of Theorem 6.1).
Let $J=\max\{\text{Pdim}(D^{2}\widetilde{\Phi}),\text{Pdim}(\widetilde{\Phi})\}$. Due to Lemma 6.10, 6.12 and Theorem 6.13, for any $M\geq J$, we have
$\mathbf{R}_{M}(D^{2}\widetilde{\Phi})\leq$ $4\delta+\frac{12}{\sqrt{M}}\int_{\delta}^{B}\sqrt{\log 2\mathcal{N}(\varepsilon,D^{2}\widetilde{\Phi},M)}\,\mathrm{d}\varepsilon$
$\leq$ $4\delta+\frac{12}{\sqrt{M}}\int_{\delta}^{B}\sqrt{\log 2\left(\frac{2eMB}{\varepsilon\text{Pdim}(D^{2}\widetilde{\Phi})}\right)^{\text{Pdim}(D^{2}\widetilde{\Phi})}}\,\mathrm{d}\varepsilon$
(69) $\leq$ $4\delta+\frac{12B}{\sqrt{M}}+12\left(\frac{\text{Pdim}(D^{2}\widetilde{\Phi})}{M}\right)^{\frac{1}{2}}\int_{\delta}^{B}\sqrt{\log\left(\frac{2eMB}{\varepsilon\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}\,\mathrm{d}\varepsilon.$
By the direct calculation for the integral, we have
$\int_{\delta}^{B}\sqrt{\log\left(\frac{2eMB}{\varepsilon\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}\,\mathrm{d}\varepsilon\leq B\sqrt{\log\left(\frac{2eMB}{\delta\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}.$
Then choosing $\delta=B\left(\frac{\text{Pdim}(D^{2}\widetilde{\Phi})}{M}\right)^{\frac{1}{2}}\leq B$, we have
(70) $\mathbf{R}_{M}(D^{2}\widetilde{\Phi})\leq 28B\left(\frac{\text{Pdim}(D^{2}\widetilde{\Phi})}{M}\right)^{\frac{1}{2}}\sqrt{\log\left(\frac{2eM}{\text{Pdim}(D^{2}\widetilde{\Phi})}\right)}.$
Therefore, due to Theorem 6.13, there is a constant $C_{4}$ independent of $L,N,M$ such as
(71) $\mathbf{R}_{M}(D^{2}\widetilde{\Phi})\leq C_{4}\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M.$
$\mathbf{R}_{M}(\widetilde{\Phi})$ can be estimate in the similar way. Due to Lemma 6.6, we have that there is a constant $C_{5}=C_{5}(B,d,C_{1},C_{2})$ such that
(72) $\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathbb{E}\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\leq C_{5}\frac{NL(\log_{2}L\log_{2}N)^{\frac{1}{2}}}{\sqrt{M}}\log M.$
6.2 Proof of Theorem 6.2
The following lemma will be used to bounded generalization error by covering numbers:
Lemma 6.16 ([16], Theorem 11.4).
Let $M\in\mathbb{N}$, and assume $\|f\|_{L^{\infty}(\Omega)}\leq B$ for some $B\geq 1$. Let $\mathcal{F}$ be a set of functions from $\mathcal{X}$ to $[-B,B]$. Then for any $20<\epsilon\leq 1/$ 2 and $\alpha,\beta>0$,
$\mathbb{P}\left\{\exists\phi\in\mathcal{F}:\left\|f-\phi\right\|_{L^{2}}^{2}-\left(\mathcal{R}_{S,0}(\phi)-\mathcal{R}_{S,0}\left(f\right)\right)\geq\epsilon\left(\alpha+\beta+\left\|f-\phi\right\|_{L^{2}}^{2}\right)\right\}$
$\leq 14\mathcal{N}\left(\frac{\beta\epsilon}{20B},\mathcal{F},M\right)\exp\left(-\frac{\epsilon^{2}(1-\epsilon)\alpha M}{214(1+\epsilon)B^{4}}\right),$
where
$\mathcal{R}_{S,0}(g)=\frac{1}{M}\sum_{i=1}^{M}|g(\boldsymbol{x}_{i})-y_{i}|^{2}$
for $\{y_{i}\}_{i=1}^{M}\subset[-B,B]$ and $\{\boldsymbol{x}_{i}\}_{i=1}^{M}\subset\Omega$.
Remark 6.17.
In the ordinary Lemma of [16], the covering number of $L^{1}$ is used to bound the inequality, which is stronger than our lemma here. Since $L^{\infty}$ is sufficient to obtain our results and to use Lemma 6.12 directly, we utilize the covering number under the $L^{\infty}$-norm here. Additionally, while their setup involves data with noise, our problem assumes noise-free data. However, the method we use can be easily generalized to accommodate noisy data.
Proposition 6.18.
Let $d,N,L,M\in\mathbb{N}_{+}$, $B,C_{1},C_{2}\in\mathbb{R}_{+}$. For any $f\in W^{n,\infty}([a,b]^{d})$ with $\|f\|_{W^{n,\infty}([a,b]^{d})}\leq B$, we have
$\mathbb{E}\mathcal{R}_{D}(\boldsymbol{\theta}_{S})\leq\frac{5136B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80BM},\widetilde{\Phi},M\right)\right)+1\right]$
(73) $+\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80dBM},D^{2}\widetilde{\Phi},M\right)\right)+1\right]+3\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{W^{2,\infty}\left([a,b]^{d}\right)}^{2},$
$\mathbb{E}$ is expected responding to $X$, and $X:=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{M}\}$ is an independent random variables set uniformly distributed on $[a,b]^{d}$.
Proof 6.19.
Due to the definition, we know that
$\mathcal{R}_{D}(\boldsymbol{\theta}_{D})=\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{H^{2}\left([a,b]^{d}\right)}^{2}.$
For the sample error, due to $\boldsymbol{\theta}_{S}$ and $\boldsymbol{\theta}_{D}$ belong to $\Theta$ almost surely, we have
$\mathbb{E}\left[\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\right]$
$\leq$ $\mathbb{E}\left[\int_{[a,b]^{d}}\left|\nabla f(\boldsymbol{x})-\nabla\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\nabla f(\boldsymbol{x}_{i})-\nabla\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]$
(74) $+\mathbb{E}\left[\int_{[a,b]^{d}}\left|f(\boldsymbol{x})-\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|f(\boldsymbol{x}_{i})-\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]$
For simplicity, we will only bound the first part; the second part can be bounded in a similar manner.
For the first sample error, we have
$\mathbb{E}\left[\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]$
$=$ $\mathbb{E}\left[\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]+\mathbb{E}\frac{(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}$
(75) $\leq$ $\mathbb{E}\left[\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\right]+\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{W^{2,\infty}\left([a,b]^{d}\right)}^{2}$
where the last inequality is due to the definition of $\boldsymbol{\theta}_{S}$.
Set
(76) $\mathbb{A}_{1}=\int_{[a,b]^{d}}\left|\Delta f(\boldsymbol{x})-\Delta\phi(\boldsymbol{x};\boldsymbol{\theta}_{S})\right|^{2}\,\mathrm{d}\boldsymbol{x}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}.$
Note
$\mathbb{P}(\mathbb{A}_{1}\geq\epsilon)$ $=\mathbb{P}\left(2\|\Delta f-\Delta\phi\|^{2}_{L^{2}([a,b]^{d})}-\frac{2(b-a)^{d}}{M}\sum_{i=1}^{M}\left|\Delta f(\boldsymbol{x}_{i})-\Delta\phi(\boldsymbol{x}_{i};\boldsymbol{\theta}_{S})\right|^{2}\geq\epsilon+\|\Delta f-\Delta\phi\|^{2}_{L^{2}([a,b]^{d})}\right)$
(77) $\leq 14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\exp\left(-\frac{\epsilon M(b-a)^{d}}{5136d^{4}B^{4}}\right),$
where the last inequality is due to Lemma 6.16.
Therefore, we have that
$\mathbb{E}\mathbb{A}_{1}$ $\leq\int_{0}^{\infty}\mathbb{P}(\mathbb{A}_{1}\geq y)\,\mathrm{d}y\leq\epsilon+\int_{\epsilon}^{\infty}\mathbb{P}(\mathbb{A}_{1}\geq y)\,\mathrm{d}y$
(78) $\leq\epsilon+\int_{\epsilon}^{\infty}14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\exp\left(-\frac{yM}{5136d^{4}B^{4}(b-a)^{d}}\right)\,\mathrm{d}y.$
By the direct calculation, we have
$\int_{\epsilon}^{\infty}14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\exp\left(-\frac{yM}{5136d^{4}B^{4}(b-a)^{d}}\right)\,\mathrm{d}y$
(79) $\leq$ $14\mathcal{N}\left(\frac{\epsilon}{80dB},D^{2}\widetilde{\Phi},M\right)\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\exp\left(-\frac{\epsilon M}{5136d^{4}B^{4}(b-a)^{d}}\right).$
Set
$\epsilon=\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80dBM},D^{2}\widetilde{\Phi},M\right)\right)\geq\frac{(b-a)^{d}}{M}$
and we have
$\mathbb{E}\mathbb{A}_{1}\leq\frac{5136B^{4}d^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80BM},D^{2}\widetilde{\Phi},M\right)\right)+1\right]$
Hence we have
$\mathbb{E}\left[\mathcal{R}_{D}(\boldsymbol{\theta}_{S})-\mathcal{R}_{S}(\boldsymbol{\theta}_{S})\right]\leq\frac{5136B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80BM},\widetilde{\Phi},M\right)\right)+1\right]$
(80) $+\frac{5136d^{4}B^{4}(b-a)^{d}}{M}\left[\log\left(14\mathcal{N}\left(\frac{(b-a)^{d}}{80dBM},D^{2}\widetilde{\Phi},M\right)\right)+1\right]+2\inf_{\phi\in\widetilde{\Phi}}\|\phi-f\|_{W^{2,\infty}\left([a,b]^{d}\right)}^{2}.$
Proof 6.20 (Proof of Theorem 6.2).
The proof can be obtained directly by combining Proposition 6.18, Lemma 6.12, and Theorem 6.13.
7 Numerical Experiment
In this section, we provide some examples to validate that DSRN can approximate smooth functions in Sobolev spaces.
7.1 General function approximation
To illustrate that the Squared ReLU activation in DSRN improves the derivative approximation, compared to neural networks using ReLU, we train DSRN and ReLU neural networks to learn the cubic function $f(x)=x^{3}$ as well as its first-order and second-order derivatives. In Figure 6, we visualize the cubic functions learned by the ReLU neural network (Figure 6(a)) and DSRN (Figure 6(d)), the first-order derivative $f^{\prime}(x)=3x^{2}$ estimated by the ReLU neural network (Figure 6(b)) and DSRN (Figure 6(e)), and the second-order derivative $f^{\prime\prime}(x)=6x$ estimated by the ReLU neural network (Figure 6(c)) and DSRN (Figure 6(f)). The comparison results demonstrate that the DSRN can approximate both the cubic function and its derivatives. In contrast, the ReLU neural network can only approximate the function up to the first-order derivative. Notably, the ReLU network fails to approximate the cubic function in terms of the second-order Sobolev norm, because its second-order derivative is zero, leading to significant errors in second-order approximation. On the other hand, the DSRN is able to learn the cubic function with the accurate first and second-order derivatives, achieving approximation in higher-order Sobolev norm.
(a) ReLU: $f(x)=x^{3}$
(b) ReLU: $f^{\prime}(x)=3x^{2}$
(c) ReLU: $f^{\prime\prime}(x)=6x$
(d) DSRN: $f(x)=x^{3}$
(e) DSRN: $f^{\prime}(x)=3x^{2}$
(f) DSRN: $f^{\prime\prime}(x)=6x$
Figure 6: Comparison of function prediction and the first-order, second-order derivatives between true value, ReLU network and DSRN.
7.2 One-dimensional Poisson equation
We solve the Poisson equation
(81) $\Delta u(x)=-\pi^{2}\sin(\pi x),\quad x\in\left[-1,1\right],$
with boundary conditions $u(-1)=u(1)=0$. The exact solution is $u(x)=\sin(\pi x)$. To be specific, we approximate the solution $u(x)$ by the neural network $\mathcal{NN}_{\boldsymbol{\theta}}(x)$ and train the neural network parameters $\boldsymbol{\theta}$ to minimize the mean squared error(MSE)
$\frac{1}{N}\sum_{i=1}^{N}\left[\left(\frac{d^{2}}{dx^{2}}\mathcal{NN}_{\boldsymbol{\theta}}(x_{i})+\pi^{2}\sin(\pi x_{i})\right)^{2}+\left(\mathcal{NN}_{\boldsymbol{\theta}}(x_{i})-y_{i}\right)^{2}\right].$
The first term comes from PINN method and the second term is used for supervised leaning with $N$ training data $\left\{(x_{i},y_{i}),i=1,\dots,N|y_{i}=\sin(\pi x_{i})\right\}$. We train 14-layer fully-connected networks to learn the solution $u(x)$. For DSRN, the activation function of the last four hidden layers is the square of ReLU, and all other activation functions are ReLU. For comparison, we also train neural networks with ReLU, and the square of ReLU as the activation function, respectively. We use the Adam optimizer with learning step $0.001$, weight decay $0.00001$ in the training.
(a) ReLU
(b) DSRN
(c) Squared ReLU
(d) DSRN
Figure 7: Comparison of prediction and the calculated first-order derivative between (a) ReLU, (b) DSRN and (c) Squared ReLU. (d) Visualization of the exact solution and the prediction, the calculated first-order, second-order derivative of DSRN.
In Figure 7, We plot the prediction of the neural network with ReLU (Figure 7(a)), DSRN (Figure 7(b)) and the square of ReLU (Figure 7(c)), respectively. We also plot the first-order derivative $\frac{d}{dx}\mathcal{NN}_{\boldsymbol{\theta}}$ using the autograd. The exact solution is $u(x)=\sin(\pi x)$, which corresponds to the ground truth in the figure. Figure 7(a) shows that the networks only with ReLU activation function is not able to learn the solution with smooth first-order derivative. However, in Figure 7(b), our method learns a solution with considerable smooth first-order derivative. On the other hand, for the neural network with squared ReLU activation functions, we observed phenomena of gradient explosion and gradient vanishing during training, attributed to the depth of networks and the square operation. Figure 7(c) shows that the network using squared ReLU suffers from gradient vanishing and therefore fails to learn in the training. These comparisons demonstrate that for deep fully connected networks, our DSRN method outperforms networks using only ReLU or squared ReLU. We further visualize the second-order derivative of the DSRN outputs in Figure 7(d). Figure 7 shows that DSRN is able to approximate the smooth function with the square of ReLU activation function only in the last several layers in the neural network. In Table 1, we list the MSE between the neural network solutions and the exact solution: DSRN achieves more accurate approximation than neural networks with only ReLU or square of ReLU. Note that the neural network with ReLU has zero-valued second-order derivative and thus a larger second-order derivative error.
Table 1: Mean squared error comparison.
ReLU DSRN Squared ReLU
Solution error $3\times 10^{-4}$ $\mathbf{1}\times\mathbf{10}^{-\mathbf{5}}$ $0.5$
Solution error + 2nd-derivative error $48.61$ $\mathbf{1.1}\times\mathbf{10}^{-\mathbf{3}}$ $49.11$
7.3 Two-dimensional Diffusion equation
We solve the diffusion equation
(82) $\frac{\partial y}{\partial t}=\frac{\partial^{2}y}{\partial x^{2}}-e^{-t}(\sin(\pi x)-\pi^{2}\sin(\pi x)),\quad x\in[-1,1],\quad t\in[0,1],$
with initial and boundary conditions $y(x,0)=\sin(\pi x),y(-1,t)=y(1,t)=0$. The exact solution is $y(x,t)=e^{-t}\sin(\pi x)$. We use a 12-layer fully-connected network $\mathcal{NN}_{\boldsymbol{\theta}}(x,t)$ to learn the solution $y(x,t)$, with the activation function of the last four hidden layers being the square of ReLU, and all other activation functions being ReLU. Similarly, the loss function is
$\frac{1}{N}\sum_{i=1}^{N}\left[\left(\frac{\partial\mathcal{NN}}{\partial t}(x_{i},t_{i})-\frac{\partial^{2}\mathcal{NN}}{\partial x^{2}}(x_{i},t_{i})+e^{-t_{i}}(sin(\pi x_{i})-\pi^{2}sin(\pi x_{i}))\right)^{2}+\left(\mathcal{NN}(x_{i},t_{i})-y_{i}\right)^{2}\right]$
where $\left\{(x_{i},t_{i},y_{i}),i=1,\dots,N|y_{i}=e^{-t_{i}}\sin(\pi x_{i})\right\}$ are training data. In the training, we use the Adam optimizer with learning step $0.001$. In Figure 8, we plot surface and contours of the DSRN output and the exact solution, respectively.
(a) DSRN output surface
(b) DSRN output contour
(c) Exact solution surface
(d) Exact solution contour
Figure 8: Comparison of surface and contour plots between DSRN predictions (a),(b) and the exact solution (c),(d).
For better comparison, we plot the contours of the DSRN output and the exact solution in one figure, see Figure 9. These comparisons demonstrate that DSRN accurately approximates the smooth solution.
Figure 9: Contour comparison: solid line for exact solution, dotted line for DSRN.
8 Conclusion
This paper introduces deep super ReLU networks (DSRNs) to approximate functions in Sobolev spaces measured by Sobolev norms, which ReLU DNNs cannot achieve. The DSRNs are constructed by adding several layers with the square of ReLU at the end of the neural networks to smooth them. While our DNNs are ReLU-based except for a few layers at the end, they retain the advantages of ReLU DNNs and are easy to train. Using other types of smooth activation functions instead of the square of ReLU may also achieve this goal, and we plan to explore these possibilities in future works. Additionally, the paper proves the optimality of DSRNs by estimating the VC-dimension of higher-order derivatives of DNNs and obtaining the generalization error in Sobolev space via the estimation of the pseudo-dimension of higher-order derivatives of DNNs. However, the method for estimating the VC-dimension of higher-order derivatives of DSRNs cannot be used for $\sigma_{2}$-NNs to obtain an optimal bound. Hence, a future research direction is to obtain optimal bounds for the VC-dimension of higher-order derivatives of $\sigma_{2}$-NNs.
acknowledgement
The work of Y. Y. are supported by National Institute of General Medical Sciences through grant 1R35GM146894. The work of H. Y. was partially supported by the US National Science Foundation under award DMS-2244988, DMS-2206333, and the Office of Naval Research Award N00014-23-1-2007. The work of Y. X. was supported by the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.